### v6

# Exact Coset Sampling forQuantum Lattice Algorithms

Abstract
We revisit the post-processing phase of Chen’s Karst-wave quantum lattice algorithm (chen2024quantum) in the Learning with Errors (LWE) parameter regime. Conditioned on a transcript $E$, the post-Step 7 coordinate state on $(\mathbb{Z}_{M})^{n}$ is supported on the shifted grid line
$\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\},$
where $\Delta=2D^{2}\mathbf{b}$ and $M=2M_{2}=2D^{2}Q$ with $Q$ odd. The amplitudes contain a quadratic chirp $\omega_{Q}^{-j^{2}}$ and an unknown run-dependent offset $\bm{v}^{\ast}(E)$. We show that Chen’s Steps 8-9 can be replaced by a single routine, Step $9^{\dagger}$: measure $\tau:=X_{1}\bmod D^{2}$, apply a $\tau$-dependent quadratic phase on the first coordinate register to remove the chirp, and then apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers. The chirp removal uses a canonical gauge promise: conditioned on $E$, the offset $v_{1}^{\ast}(E)\bmod M_{2}$ equals the centered lift $\operatorname{ctr}(\tau)$. Under explicit front-end conditions (AC1-AC5), a Fourier sample $\bm{u}\in\mathbb{Z}_{M}^{n}$ satisfies the resonance $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ with probability $1-\operatorname{negl}(n)$, and conditioned on resonance the reduced sample $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane. This yields exact recovery of $\mathbf{b}\bmod Q$ and hence the embedded LWE secret.
1 Introduction
[…]
for an outcome-dependent amplitude profile $\alpha_{E}(j,\bm{k})$ and an index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$ that comes from the grid-state construction. When it is clear from context, we suppress the transcript index $E$. The phase factor $i^{\|\bm{k}\|^{2}}$ comes from Chen’s grid-state preparation procedure. We view the coordinate values as elements of $(\mathbb{Z}_{M})^{n}$ throughout. Accordingly, $\mathbf{X}(j):=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ is interpreted componentwise modulo $M$ (as in Chen’s post-Step 7 expression), and $\bm{k}$ records the $M_{2}$-shift term $M_{2}\bm{k}$. Reducing the coordinate registers modulo $M_{2}$ eliminates this shift and yields the affine congruence $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}\pmod{M_{2}}$ used in Additional Condition AC1.
[…]
Additional Condition AC1 records the affine relation $\mathbf{X}(j)=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ for the post-Step 7 state. Additional Condition AC4 fixes a canonical gauge for the first-coordinate offset: the preparation circuit is compiled (by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$) so that measuring $\tau:=X_{1}\bmod D^{2}$ determines $v_{1}^{\ast}(E)\bmod Q$ for that run. Under Additional Condition AC1, the first coordinate satisfies the precise congruence
$X_{1}(j)\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}}$
for every $j$ in the effective window. Write $\Delta:=2D^{2}\mathbf{b}$ and recall $M=2M_{2}=2D^{2}Q$. Since $D^{2},Q\mid M$, reducing any physical basis value $x\in\mathbb{Z}_{M}$ modulo $D^{2}$ or $Q$ is unambiguous. In particular,
$X_{1}\bmod D^{2}\equiv v_{1}^{\ast}\bmod D^{2},\qquad(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$,
where under the canonical gauge (AC4) the post-processing computes $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from $\tau=X_{1}\bmod D^{2}$.
In Chen’s post-processing one ultimately outputs non-zero Fourier samples in $\mathbb{Z}_{M_{2}}^{n}$ (with $M_{2}=M/2$) whose reduction modulo $Q=M/(2D^{2})$ satisfies a homogeneous dual relation. In our direct-QFT variant we instead measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ and store only the reduction $\bm{u}_{Q}:=\bm{u}\bmod Q$. In either view, the goal is to obtain non-zero vectors satisfying
$\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, (1.3)
where $Q:=\frac{M}{2D^{2}}$ is an odd integer. (In chen2024quantum the parameters are additionally arranged so that $Q$ has a convenient factorization, but Step $9^{\dagger}$ does not use it.) The short vector $\mathbf{b}=(b_{1},\dots,b_{n})\in\mathbb{Z}^{n}$ stores the hidden data and satisfies the normalization $b_{1}=-1$. The offset $\bm{v}^{\ast}$ is unknown and can change from run to run. The algorithm then collects $O(n)$ such samples and solves the resulting linear system modulo $Q$ to recover $\mathbf{b}\bmod Q$ and hence the hidden data vector.
Chen’s published Steps 8-9 enforce relations of the form Equation (1.3) via additional offset-handling and a domain-extension mechanism that is tailored to a particular parameterization. Step $9^{\dagger}$ provides an alternative post-processing designed for an access model in which the affine offset $\bm{v}^{\ast}(E)$ may vary across runs and no special factorization of $Q$ is assumed. It uses only the affine structure $\mathbf{X}(j)=j\Delta+\bm{v}^{\ast}$ with $\Delta=2D^{2}\mathbf{b}$, the normalization $b_{1}=-1$, a run-local chirp cancellation acting on $X_{1}$ (Additional Condition AC4), and a direct $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ on the coordinate registers. Under Additional Condition AC5, the resulting outcomes satisfy $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ except with negligible failure probability.
When we instantiate the front end with the $q$-ary lattice $L_{q}^{\perp}(\mathbf{A})$ from chen2024quantum, the new Step $9^{\dagger}$ recovers the planted shortest vector $\mathbf{b}=[-1,2p_{1}s^{\top},2p_{1}e^{\top}]^{\top}$ and hence the LWE secret $s$ and error $e$ in the chosen-secret regime considered there. Combined with the classical reductions of chen2024quantum, this yields a quantum algorithm for standard LWE in the same parameter regime, conditional on Additional Conditions AC1-AC5.
2 Background and Access Model
Notations.
Let $q\in\mathbb{N}$ and write $\mathbb{Z}_{q}=\mathbb{Z}/q\mathbb{Z}$ with representatives in $(-\tfrac{q}{2},\tfrac{q}{2}]$. Vectors are written in bold. We use $\langle\cdot,\cdot\rangle$ for the inner product. For $T\in\mathbb{N}$ we write $\omega_{T}:=\exp(2\pi i/T)$. The physical coordinate registers live in $\mathbb{Z}_{M}^{n}$, where
$M:=2M_{2}=2D^{2}Q,\qquad M_{2}:=D^{2}Q.$
We will repeatedly reduce physical basis values $x\in\mathbb{Z}_{M}$ modulo $M_{2}$, $D^{2}$, or $Q$. Since $D^{2},Q\mid M_{2}\mid M$, the canonical reduction maps
$\mathbb{Z}_{M}\to\mathbb{Z}_{M_{2}},\qquad\mathbb{Z}_{M}\to\mathbb{Z}_{D^{2}},\qquad\mathbb{Z}_{M}\to\mathbb{Z}_{Q}$
are well-defined ring homomorphisms.
Modular inner products.
For $\bm{a}\in\mathbb{Z}^{n}$ and $\bm{u}\in(\mathbb{Z}_{M})^{n}$ we write
$\langle\bm{a},\bm{u}\rangle\bmod Q\ :=\ \sum_{i=1}^{n}(a_{i}\bmod Q)\cdot(u_{i}\bmod Q)\ \in\ \mathbb{Z}_{Q}.$
When we write $\langle\bm{a},\bm{u}\rangle\equiv 0\pmod{Q}$ this is the intended meaning (i.e., the dot product after coordinate-wise reduction $\mathbb{Z}_{M}\to\mathbb{Z}_{Q}$).
Representative conventions.
[…]
We write
$0\!\mid\!\mathbb{Z}^{n-1}:=\{0\}\times\mathbb{Z}^{n-1}.$
The physical coordinate block in Chen’s Step 7 lives in $\mathbb{Z}_{M}$ with $M=2M_{2}$. When this helps to avoid confusion, we write the reduction modulo $M$ with a subscript M. For a coordinate slice we write $\bm{x}_{[2..n]}:=(x_{2},\dots,x_{n})$. Set $\Delta:=2D^{2}\,\mathbf{b}\in\mathbb{Z}^{n}$.
We use $\operatorname{poly}(\cdot)$ for an unspecified fixed polynomial and $\operatorname{negl}(n)$ for a negligible function of $n$.
Idealized unitaries vs. circuit precision.
For clarity we treat $\mathrm{QFT}_{\mathbb{Z}_{M}}$ and the controlled root-of-unity phases used in the chirp correction (AC4) as exact unitaries. In a standard uniform gate set these operations can be implemented to precision $2^{-\operatorname{poly}(n)}$, which perturbs all output distributions by at most $\operatorname{negl}(n)$ in total variation distance. We suppress this approximation issue throughout, since it does not affect any of the modular-algebraic arguments.
Finally, we write $\operatorname{ctr}:\mathbb{Z}_{D^{2}}\to\mathbb{Z}$ for the centered lift: for $\tau\in\mathbb{Z}_{D^{2}}$, $\operatorname{ctr}(\tau)$ is the unique integer in $(-D^{2}/2,D^{2}/2]$ that is congruent to $\tau$ modulo $D^{2}$. We use $\operatorname{ctr}(\tau)$ as a canonical integer representative when converting $\tau=X_{1}\bmod D^{2}$ into arithmetic in $\mathbb{Z}_{Q}$.
At the frontier right before Step 8, and for fixed outcomes $E$, Lemma 2.8 gives $\mathbf{X}(j)\equiv j\Delta+\bm{v}^{\ast}\pmod{M_{2}}$ for every $j$ in the effective window.
Parameter identification from chen2024quantum.
In Chen’s notation, $M=2(t^{2}{+}u^{2})$ and $x=D\mathbf{b}$, where $\mathbf{b}$ is the short vector from the LWE-to-lattice reduction (Eq. (12) in chen2024quantum). We follow this identification. Chen’s Step 9 defines a relabelled vector $\mathbf{b}^{\ast}$ (Eq. (39) in that paper). Our post-processing never uses $\mathbf{b}^{\ast}$. We keep all formulas in terms of $\mathbf{b}$ instead.
Under Chen’s parameter constraints the quantity $M/(2D^{2})$ is an odd integer. We write
$Q:=\frac{M}{2D^{2}},\qquad M_{2}:=\frac{M}{2}=D^{2}Q.$
In Chen’s concrete instantiation $Q$ happens to admit a factorisation into pairwise coprime odd factors; the new Step $9^{\dagger}$ does not use this extra structure and only relies on $Q$ being odd.
Chen uses the letter $P$ for a large loop modulus with $P\approx M^{2}/2$. We keep this convention and reserve $P$ only for that modulus.
Index domain vs. first coordinate.
The preparation uses a loop index $j$. This index is the discrete time variable of the Karst wave and appears in chen2024quantum as a summation index. It runs over the large modulus $P$. We denote the register by $J\in\mathbb{Z}_{P}$.
The first coordinate satisfies
$X_{1}\equiv 2D^{2}j\,b_{1}+v_{1}^{\ast}\pmod{M_{2}}$.
In Chen’s setup, $v_{1}^{\ast}$ depends on the measurement outcomes $(y^{\prime},z^{\prime},h^{\star})$. The experiment cannot be reset in a controlled way to obtain the same outcomes and the same $v^{\ast}$. So we cannot query the state preparation on chosen basis inputs $j$ and then compute $\mathbf{X}(j)$ either classically or into an extra register. We must work with the single quantum state $|\varphi_{7}\rangle$ that a successful run outputs.
Parameter re-tuning.
As in chen2024quantum we run a classical outer loop over polynomially many candidates for $\|\mathbf{b}\|^{2}$ and, for each guess, set
$Q=(c{+}1)\|\mathbf{b}\|^{2},\qquad M=2D^{2}Q,\qquad M_{2}=D^{2}Q.$
The scaling parameter $D$ can then be chosen (odd, $\gcd(D,Q)=1$, and $\operatorname{poly}(n)$-bounded) to satisfy the geometric separation requirement of the Karst-wave front end (Condition C.6 of chen2024quantum). The only additional analytic requirement used by Step $9^{\dagger}$ is the one-dimensional spectral concentration property recorded in Additional Condition AC5, which is achieved in Chen’s complex-Gaussian regime once the loop-envelope width satisfies $\sigma_{J}\gtrsim Q\log n$ (Remark 2.5). All subsequent correctness statements are conditional on the existence of a (re-tuned) front-end configuration satisfying Definition 2.1 and Additional Conditions AC1-AC5.
Definition 2.1 (Re-tuned Karst-wave configuration).
Take LWE parameters $(\ell,m,q,\beta)$ in the regime of chen2024quantum and set $n:=1+\ell+m$. Let $L_{q}^{\perp}(\mathbf{A})$, the planted shortest vector $\mathbf{b}$, and the integers $(p_{\eta})_{\eta}$ be as in Section 3.2 of chen2024quantum. A re-tuned Karst-wave configuration is any choice of front-end parameters
$(c,D,r,s,t,u,P,M,Q)$
that satisfies the following properties.
  1. (R1)
    The integers $p_{1},\dots,p_{\kappa}$ are odd and pairwise coprime. They control the chosen-secret embedding and the parsing of $\mathbf{b}$ in the LWE extraction step, but they do not impose any algebraic constraint on $Q$. Fix an integer $c=c(n)\in 4\mathbb{Z}$ (typically $c=\Theta(\log^{6}n)$ in Chen-style instantiations) and set
    $Q:=(c{+}1)\|\mathbf{b}\|^{2}.$
    In the chosen-secret LWE embedding of (chen2024quantum), $\mathbf{b}=[-1,2p_{1}s^{\top},2p_{1}e^{\top}]^{\top}$ has one odd coordinate and all remaining coordinates even, hence $\|\mathbf{b}\|^{2}\equiv 1\pmod{4}$. Since $c\in 4\mathbb{Z}$, this makes $Q=(c{+}1)\|\mathbf{b}\|^{2}$ automatically odd. We also require that $Q$ satisfies
    $\|\mathbf{b}\|_{\infty}\leq 2p_{1}\beta\log n<Q/2,$
    so that $\mathbf{b}\bmod Q$ uniquely determines $\mathbf{b}$ by centered lifting.
  2. (R2)
    Choose an odd scaling parameter $D$ that is coprime to $Q$ and polynomially bounded in $n$ (for concreteness, one may take the smallest odd integer at least $32\log^{2}n$ with $\gcd(D,Q)=1$). Define
    $u^{2}:=\|D\mathbf{b}\|^{2}=D^{2}\|\mathbf{b}\|^{2},\qquad t^{2}:=c\,u^{2},$
    and set the Karst-wave moduli
    $M:=2(t^{2}+u^{2})=2D^{2}Q,\qquad M_{2}:=M/2=D^{2}Q,\qquad P:=M(t^{2}+u^{2})=\frac{M^{2}}{2}.$
    Then $2D^{2}$ is a unit modulo $Q$.
  3. (R3)
    Choose $r,s>0$ so that Chen’s Karst-wave constraints C.5-C.7 hold and the resulting geometric width parameter
    $\sigma_{\mathrm{geom}}:=\frac{rs^{2}t}{u\sqrt{r^{4}+s^{4}}}$
    obeys $2\log n<\sigma_{\mathrm{geom}}<D/(4\log n)$. Chen’s analysis shows that such choices exist with all parameters bounded by $\operatorname{poly}(n)$; we fix one such choice.
  4. (R4)
    For the resulting complex-Gaussian front end, let $\sigma_{J}$ denote the standard deviation of the one-dimensional envelope $\alpha_{E}(j)$ along the loop index $j$ (after marginalizing over the grid index and the workspace and conditioning on a transcript $E$). We restrict to parameter choices satisfying
    $\sigma_{J}\geq C_{\mathrm{spec}}\,Q\log n$
    for some fixed absolute constant $C_{\mathrm{spec}}>0$. A lower bound of this form keeps the spectral leakage in Additional Condition AC5 bounded by $\exp(-\Omega(\log^{2}n))$.
When we refer to Chen’s front end with re-tuned parameters below, we mean an outcome-conditioned circuit family compiled with a fixed choice of parameters that satisfies (R1)-(R4).
2.1 Front-end properties and additional conditions
We now record the structural properties and spectral conditions that the re-tuned front end must satisfy. These conditions allow the LWE solver in Step $9^{\dagger}$ to work in the exact sampling regime.
AC1.
After modulus splitting and center extraction (Steps 5-7 of chen2024quantum) and just before the Step 8 gadget, conditioned on any fixed transcript $E$ the post-Step 7 coordinate block is supported on
$\bigl\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\bigr\},$
for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields the affine congruence
$\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}(E)\pmod{M_{2}}$.
This matches the offset-coherence invariant of Chen’s deferred-measurement front end (Lemma 2.8). In particular, since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$ and $b_{1}=-1$, the first coordinate satisfies
$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}\pmod{M_{2}},\qquad X_{1}\bmod D^{2}\equiv v_{1}^{\ast}\bmod D^{2}.$
AC2.
We assume throughout that $Q$ is an odd integer. The post-processing in Step $9^{\dagger}$ works for an arbitrary composite modulus $Q$; whenever we invert a residue class we only do so modulo a modulus that is guaranteed to be coprime to that residue (in particular, we only invert $2D^{2}$ modulo $Q$, which is valid by AC2-AC3). In particular, we do not require $Q$ to factor in any prescribed way, and the integers $(p_{\eta})_{\eta}$ from the LWE-to-lattice reduction play no role in the algebraic structure of $Q$.
AC3.
We keep the scaling regime of chen2024quantum. The parameter $D$ is odd, coprime to $Q$, and we define
$M:=2D^{2}Q,\qquad M_{2}:=M/2=D^{2}Q.$
Since $Q$ is odd (AC2) and $\gcd(D,Q)=1$, we have $\gcd(2D^{2},Q)=1$, so $2D^{2}$ is a unit modulo $Q$. All divisions by $D^{2}$ or $2D^{2}$ that appear in the post-processing should be read as multiplications by the corresponding inverses in $\mathbb{Z}_{Q}$; we never perform integer division by $D^{2}$ or $2D^{2}$ in $\mathbb{Z}$. For the first coordinate, this gives
$X_{1}\equiv 2D^{2}jb_{1}+v_{1}^{\ast}\equiv-2D^{2}j+v_{1}^{\ast}\pmod{M_{2}}$,
so that reduction modulo $D^{2}$ kills the $j$-dependent term:
$X_{1}\bmod D^{2}\equiv v_{1}^{\ast}\bmod D^{2}.$
AC4.
The post-Step 7 amplitudes contain the quadratic chirp $\omega_{Q}^{-j^{2}}$ in the loop index $j$. We cancel it with a diagonal unitary acting only on the first coordinate register $X_{1}$.
Fix a transcript $E$. By AC1 and $b_{1}=-1$, for every $j$ in the effective window,
$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}}$.
Because $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$, reducing modulo $D^{2}$ removes the $j$-term on every branch:
$\tau:=X_{1}\bmod D^{2}\equiv v_{1}^{\ast}(E)\bmod D^{2}.$
Thus $\tau$ is deterministic conditioned on $E$; measuring it does not disturb the coherent superposition over $j$ and $\bm{k}$. Let $\bar{v}_{1}:=\operatorname{ctr}(\tau)\in(-D^{2}/2,D^{2}/2]$ and set $v_{1,Q}:=\bar{v}_{1}\bmod Q\in\mathbb{Z}_{Q}$.
Under the canonical-gauge requirement below, this $v_{1,Q}$ equals $v_{1}^{\ast}(E)\bmod Q$ for that run.
Canonical gauge (required). Measuring $\tau$ determines $v_{1}^{\ast}(E)\bmod D^{2}$ but leaves an ambiguity by multiples of $D^{2}$ inside $\mathbb{Z}_{M_{2}}$. Equivalently, there is a unique $\lambda(E)\in\mathbb{Z}_{Q}$ such that
$v_{1}^{\ast}(E)\equiv\bar{v}_{1}+\lambda(E)D^{2}\pmod{M_{2}}$.
Any correction depending only on $\tau$ cancels $\omega_{Q}^{-j^{2}}$ only up to a residual run-dependent linear phase $\omega_{Q}^{-\lambda(E)j}$ (up to a global constant), which shifts the resonance class $\langle\mathbf{b},\bm{u}\rangle\bmod Q$ and prevents aggregating samples across invocations. We therefore assume a canonical gauge in which $\lambda(E)=0$ for every transcript $E$, i.e.,
$v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$.
When the preparation circuit is known, this can be enforced by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$, provided $E$ allows reversible computation of $\lambda(E)$; see Proposition 2.3.
Define the classical inverse $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$, which exists by AC2-AC3. For each basis value $x\in\mathbb{Z}_{M}$ of $X_{1}$, write $x_{Q}:=x\bmod Q$ and define
$j_{x}:=\mathrm{inv}\cdot\bigl((x_{Q}-v_{1,Q})\bmod Q\bigr)\in\mathbb{Z}_{Q}.$
For a branch with loop label $j$ we have $(x_{Q}-v_{1,Q})\equiv-2D^{2}j\pmod{Q}$, hence $j_{x}\equiv j\pmod{Q}$. We apply the diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$:
[…]
AC5.
For the rephased chirp-free state of AC4 and a fixed transcript $E$, define for $\bm{u}\in\mathbb{Z}_{M}^{n}$
$S_{E}(\bm{u})\,:=\,\sum_{j\in\mathbb{Z}}\sum_{\bm{k}\in\mathcal{K}}\alpha_{E}(j,\bm{k})\,i^{\|\bm{k}\|^{2}}\,\omega_{M}^{\,j\langle\bm{u},\Delta\rangle+M_{2}\langle\bm{u},\bm{k}\rangle}.$
[…]
We view the loop index $j\in\mathbb{Z}_{P}$ as embedded in $\mathbb{Z}$. We extend $\alpha_{E}(j,\bm{k})$ by zero outside an interval $\{|j|\leq J_{\max}\}$. The effective window satisfies $J_{\max}=\operatorname{poly}(Q\log n)\ll P$ in the Karst-wave regime below; in particular we assume $J_{\max}<P/2$ so that the embedding of $j\in\mathbb{Z}_{P}$ into integers is unambiguous on the support of the state. The infinite sum above then matches, up to negligible error, the finite sum that appears in Chen’s state.
As in Chen’s Karst-wave front end, we assume that conditioned on $E$ the amplitudes factor across the loop index and the grid label:
$\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\,\beta_{E}(\bm{k}).$
Define the one-dimensional Fourier transform of the loop envelope
$\widehat{\alpha}_{E}(\theta):=\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\theta}.$
A transcript $E$ is called good if the off-resonant Fourier mass on the $Q$-point grid $\{t/Q:t\in\mathbb{Z}_{Q}\}$ is negligible:
$\sum_{t\in\mathbb{Z}_{Q}\setminus\{0\}}\bigl|\widehat{\alpha}_{E}(t/Q)\bigr|^{2}\leq\operatorname{negl}(n)\cdot\sum_{t\in\mathbb{Z}_{Q}}\bigl|\widehat{\alpha}_{E}(t/Q)\bigr|^{2}.$ (2.1)
We assume that a transcript drawn from the front end is good except with probability $\operatorname{negl}(n)$ over the front-end randomness.
Reliability of the additional conditions.
Additional Conditions AC2-AC3 are enforced directly by Definition 2.1. Additional Condition AC1 is the standard offset-coherence invariant of Chen’s deferred-measurement front end (Lemma 2.8). Additional Condition AC4 is the only genuinely new access-model promise used by Step $9^{\dagger}$: it replaces Chen’s Step 8 offset-extraction gadget by a gauge-fixing guarantee (or coherent access to the gauge parameter) sufficient to cancel the quadratic chirp using only $X_{1}$. It is enforceable by a compile-time basis translation on the first coordinate whenever the preparation circuit supports reversible computation of the gauge parameter $\lambda(E)$ (Proposition 2.3). Finally, Additional Condition AC5 follows from the complex discrete-Gaussian Karst-wave envelope once $\sigma_{J}\geq C_{\mathrm{spec}}Q\log n$ (Proposition 2.6).
Lemma 2.2 (Uniformity modulo $Q$ on the dual hyperplane).
[…]
$R:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\langle\mathbf{b},\bm{u}\rangle\equiv 0\!\!\!\pmod{Q}\},\qquad H:=\{\bm{v}\in\mathbb{Z}_{Q}^{n}:\langle\mathbf{b},\bm{v}\rangle\equiv 0\!\!\!\pmod{Q}\}$.
Conditioned on $R$, the reduced outcome $\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ is exactly uniform on $H$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=1/|H|=Q^{-(n-1)}$.
Proof.
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals
$\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}\cdot\sum_{j\in\mathbb{Z}}\sum_{\bm{k}\in\mathcal{K}}\alpha_{E}(j,\bm{k})\,i^{\|\bm{k}\|^{2}}\,\omega_{M}^{\,j\langle\bm{u},\Delta\rangle+M_{2}\langle\bm{u},\bm{k}\rangle},$
so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
Because $M_{2}=M/2$, we have $\omega_{M}^{M_{2}\langle\bm{u},\bm{k}\rangle}=(-1)^{\langle\bm{u},\bm{k}\rangle}$, which depends only on the parity pattern of $\bm{u}_{[2..n]}$ since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$. On the event $R$ we have $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, hence
$\langle\bm{u},\Delta\rangle=2D^{2}\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{2D^{2}Q}=0\pmod{M}$,
and therefore $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=1$ for all $j$. Therefore, conditioned on $R$, the outcome probabilities depend on $\bm{u}$ only through the parity pattern of $\bm{u}_{[2..n]}$.
Now fix any $\bm{v}\in H$. The fibre $F_{\bm{v}}:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\bm{u}\bmod Q=\bm{v}\}$ consists of the vectors $\bm{u}=\bm{v}+Q\bm{t}$ with $\bm{t}\in\mathbb{Z}_{M/Q}^{n}=\mathbb{Z}_{2D^{2}}^{n}$. Since $Q$ is odd, as $\bm{t}$ ranges over $\mathbb{Z}_{2D^{2}}^{n}$ the parity pattern of $(\bm{v}+Q\bm{t})_{[2..n]}$ is exactly uniform over $\{0,1\}^{n-1}$, independently of $\bm{v}$. Hence each parity pattern occurs in $F_{\bm{v}}$ with the same multiplicity for every $\bm{v}\in H$, and since the conditional probabilities depend only on this parity pattern, $\mu_{E}[F_{\bm{v}}\mid R]$ is constant over $\bm{v}\in H$. This proves uniformity of $\bm{u}\bmod Q$ on $H$ conditioned on $R$. Finally, because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, $|H|=Q^{n-1}$, so $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=Q^{-(n-1)}$. ∎
Proposition 2.3 (Canonical-gauge compilation given coherent transcript access).
Assume AC1 and suppose the state-preparation procedure is given as an explicit circuit that admits a deferred-measurement implementation producing a coherent transcript register $E$ from which $v_{1}^{\ast}(E)\bmod M_{2}$ (equivalently, the gauge parameter $\lambda(E)$ defined in AC4) can be computed reversibly. Then one can compile the preparation by appending an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, the resulting offset satisfies the canonical-gauge requirement $v_{1}^{\ast}(E)\equiv\operatorname{ctr}(X_{1}\bmod D^{2})\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$). If the preparation is a black-box state oracle that does not expose such coherent access to $\lambda(E)$, the canonical-gauge condition is an additional oracle promise and is not implied by AC1 alone.
Proof.
[…]
Lemma 2.4 (Center-referenced chirp cancellation).
Assume AC1-AC4. Then the procedure in AC4 defines, for each measured $\tau$, a unitary $U_{\mathrm{corr}}^{(\tau)}$ on the first coordinate register such that $|\varphi_{7}\rangle:=U_{\mathrm{corr}}|\widetilde{\varphi}_{7}\rangle$ has the form (1.2). The quadratic chirp $e^{-2\pi ij^{2}/Q}$ in the loop index is removed on every branch.
Proof.
Under AC1, we have $X_{1}\equiv-2D^{2}j+v_{1}^{\ast}\pmod{M_{2}}$ with $v_{1}^{\ast}$ independent of $j$. Conditioned on the transcript $E$, the measurement of $\tau:=X_{1}\bmod D^{2}$ is deterministic and yields $\tau\equiv v_{1}^{\ast}\bmod D^{2}$ on every branch. By the gauge choice in AC4, with $\bar{v}_{1}=\operatorname{ctr}(\tau)$ and $v_{1,Q}=\bar{v}_{1}\bmod Q$, we have $(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$ on every branch. The reversible computation of $j_{x}$ in AC4 is well-defined modulo $Q$ by AC2-AC3. For $x=X_{1}$ on a branch with loop label $j$ we have $j_{x}\equiv j\pmod{Q}$. The unitary $U_{\mathrm{corr}}$ multiplies that branch by $\exp(2\pi ij^{2}/Q)$. It cancels the original chirp $e^{-2\pi ij^{2}/Q}$. The map is diagonal in the computational basis of $X_{1}$. Hence it is unitary. ∎
Remark 2.5 (Justification of Additional Condition AC5).
Additional Condition AC5 isolates the only additional front-end property used by Step $9^{\dagger}$: spectral concentration of the one-dimensional Karst-wave envelope at the resonant frequencies $t/Q$ with $t\equiv 0\pmod{Q}$.
In Chen’s Karst-wave front end, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ (see Eq. (35) in chen2024quantum). After chirp cancellation (Additional Condition AC4), the Fourier amplitude at $\bm{u}\in\mathbb{Z}_{M}^{n}$ factorizes as
$S_{E}(\bm{u})=\widehat{\alpha}_{E}(t(\bm{u})/Q)\cdot G_{E}(\bm{u}),$
where $G_{E}(\bm{u})$ collects the grid contribution. Here we write
$t(\bm{u})\ :=\ \langle\mathbf{b},\bm{u}\rangle\bmod Q\ \in\ \mathbb{Z}_{Q},$
and identify $t(\bm{u})$ with its standard representative in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$ on $[0,1)$. Taking the envelope width $\sigma_{J}\gtrsim Q\log n$ yields $|\widehat{\alpha}_{E}(t/Q)|\leq\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|$ for all integers $t\not\equiv 0\pmod{Q}$ by standard Poisson-summation estimates. Since $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$, this implies the leakage bound (2.1).
Remark on linear phases. A time shift $j\mapsto j-j_{0}$ multiplies $\widehat{\alpha}_{E}(\theta)$ by a unit-modulus factor and does not affect its magnitude. In contrast, a nontrivial linear phase $\alpha_{E}(j)\mapsto\alpha_{E}(j)e^{2\pi i\nu j}$ shifts the spectrum $\widehat{\alpha}_{E}(\theta)\mapsto\widehat{\alpha}_{E}(\theta+\nu)$ and can move the dominant residue class away from $t=0$ on the grid $\{t/Q\}$. In our setting, such a run-dependent shift is exactly the residual $\omega_{Q}^{-\lambda(E)j}$ phase arising from a gauge mismatch in AC4; the canonical-gauge promise rules it out.
No additional near-uniformity assumption is needed for the reduced samples: Lemma 2.2 shows that conditioned on $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, the reduction $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane $H$, and $\Pr[\bm{u}\bmod Q=\bm{0}\mid\langle\mathbf{b},\bm{u}\rangle\equiv 0]=1/|H|=Q^{-(n-1)}$.
Proposition 2.6 (Spectral concentration for the re-tuned Karst wave).
For the complex-Gaussian Karst-wave front end of chen2024quantum (where, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$), instantiated with any re-tuned configuration that satisfies Definition 2.1, Additional Condition AC5 holds (i.e., the leakage bound (2.1) fails only with probability $\operatorname{negl}(n)$ over the front-end transcript).
Proof.
Under Chen’s Karst-wave analysis, conditioned on a transcript $E$ the loop envelope $\alpha_{E}(j)$ is (up to a global phase, an integer shift in $j$, and negligible truncation) a discrete Gaussian of width $\sigma_{J}$, and the factorization $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ holds. An integer shift in $j$ does not affect $|\widehat{\alpha}_{E}(\cdot)|$. Any additional run-dependent linear phase on the $1/Q$ grid would shift the dominant residue class away from $t=0$; in our setting this is precisely the gauge-mismatch phenomenon discussed and ruled out in AC4. Thus it suffices to bound the centered discrete-Gaussian prototype $g_{\sigma}(j):=\exp(-\pi j^{2}/\sigma^{2}).$ Its Fourier series satisfies the Poisson-summation identity
$\sum_{j\in\mathbb{Z}}g_{\sigma}(j)\,e^{2\pi ij\theta}=\sigma\sum_{m\in\mathbb{Z}}\exp\bigl(-\pi\sigma^{2}(\theta-m)^{2}\bigr).$
Evaluating at $\theta=t/Q$, for any integer $t\not\equiv 0\pmod{Q}$ let $m^{\star}\in\mathbb{Z}$ be a nearest integer to $t/Q$ (so $m^{\star}\in\{0,1\}$). Then $|t/Q-m^{\star}|\geq 1/Q$, hence the dominant Gaussian term in the Poisson sum is at most $\sigma\exp(-\pi\sigma^{2}/Q^{2})$, and all other terms are smaller by an additional factor $\exp(-\Omega(\sigma^{2}))$. Consequently, for some absolute constant $c_{0}>0$,
$|\widehat{\alpha}_{E}(t/Q)|\leq|\widehat{\alpha}_{E}(0)|\cdot\exp\Bigl(-c_{0}\cdot(\sigma_{J}/Q)^{2}\Bigr)\quad\text{for all }t\not\equiv 0\pmod{Q}$,
up to negligible truncation error from the effective window. With $\sigma_{J}\geq C_{\mathrm{spec}}Q\log n$ (Definition 2.1), this gives $|\widehat{\alpha}_{E}(t/Q)|\leq|\widehat{\alpha}_{E}(0)|\exp(-\Omega(\log^{2}n)).$ Since $Q=\operatorname{poly}(n)$ in the re-tuned regime, summing the squared bound over the (Q1)(Q-1) nonzero residues yields
$\sum_{t\in\mathbb{Z}_{Q}\setminus\{0\}}|\widehat{\alpha}_{E}(t/Q)|^{2}\leq(Q-1)\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|^{2}.$
Using $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$ proves (2.1). ∎
2.2 Access model
Definition 2.7 (Access model).
We assume (circuit-level) quantum access to an outcome-conditioned state-preparation procedure that, on each invocation, outputs one copy of the post-Step 7 state $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$. We additionally assume a canonical-gauge implementation as in Additional Condition AC4: inside the preparation circuit the first coordinate is translated by an $E$-dependent multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, writing $\tau:=X_{1}\bmod D^{2}$ and $\bar{v}_{1}:=\operatorname{ctr}(\tau)$, the offset satisfies $v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$ in Additional Condition AC4). Consequently, measuring $\tau$ allows the post-processing to compute $v_{1,Q}=\bar{v}_{1}\bmod Q$ and the chirp-correction phase for that run.
For each invocation there exists an unobserved classical transcript $E$ such that, conditioned on $E$, the reduced state on $X$ satisfies Additional Condition AC1 with offset $\bm{v}^{\ast}(E)$ and amplitudes $\alpha_{E}(j,\bm{k})$. Across invocations the transcript is re-sampled, so the offset $\bm{v}^{\ast}(E)$ may change from run to run and cannot be forced to repeat.
The post-processing is allowed to apply any quantum circuit of size $\operatorname{poly}(n)$ to the output registers, including arithmetic in $\mathbb{Z}_{M}$ and $\mathbb{Z}_{Q}$, intermediate measurements, and $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$. We do not assume the ability to query the preparation on chosen basis inputs $j$, to classically compute the full offset $\bm{v}^{\ast}(E)$, or to obtain two copies of $|\widetilde{\varphi}_{7}\rangle$ with the same offset.
In this model, an algorithm acts on the superposition directly. The offset $\bm{v}^{\ast}(E)$ depends on the run-specific outcomes $E$ and behaves like a one-time pad on the coordinate values in the computational basis. Step $9^{\dagger}$ shows that this offset turns into a harmless phase pattern after a Fourier transform. We can then sample from the dual lattice without ever learning $\Delta$ or $\bm{v}^{\ast}$ as explicit classical data.
Lemma 2.8 (Offset coherence).
In the deferred-measurement unitary of chen2024quantum, take the program point just before Step 8 and after modulus splitting and center extraction (Steps 5-7). For each fixed transcript $E$, the post-Step 7 coordinate block in $(\mathbb{Z}_{M})^{n}$ is supported on
$\bigl\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\bigr\},$
for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$. Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields
$\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}(E)\pmod{M_{2}}$.
Proof.
This is immediate from Chen’s post-Step 7 computational-basis form (see the excerpted Eq. (14)): the basis values are $2Dj\,\bm{x}+\bm{v}^{\prime}+\tfrac{M}{2}\bm{k}\bmod M$ with $\bm{k}\in 0\!\mid\!\mathbb{Z}^{n-1}$ and $\bm{x}=D\mathbf{b}$. Substituting $\bm{x}=D\mathbf{b}$ gives $2Dj\,\bm{x}=2D^{2}j\,\mathbf{b}=j\Delta$ with $\Delta=2D^{2}\mathbf{b}$, and writing $\bm{v}^{\ast}(E):=\bm{v}^{\prime}\bmod M$ yields the claimed support in $(\mathbb{Z}_{M})^{n}$. Reducing modulo $M_{2}=M/2$ eliminates the $M_{2}\bm{k}$ term, giving $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}(E)\pmod{M_{2}}$. ∎
3 The new Step $9^{\dagger}$
Algorithm 1 Step $9^{\dagger}$
1:Input: A lattice instance $L_{q}^{\perp}(\mathbf{A})$ arising from a chosen-secret LWE instance as in chen2024quantum (i.e., $\mathbf{A}=[2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$) together with public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
2:Oracle: State-preparation access to $|\widetilde{\varphi}_{7}\rangle$ as in Definition 2.7.
3:Parameters: Re-tuned Karst-wave configuration $(c,D,r,s,t,u,P,M,Q)$ (see Remark 3.1).
4:Output: Either the LWE secret $s$ and error $e$, or fail (meaning the current norm guess is rejected).
5:Fix absolute constants $C_{\mathrm{rank}}>1$, $C_{\mathrm{try}}\geq 1$, and $C_{\mathrm{samp}}\geq 1$.
6:Set the target sample count $N\leftarrow\lceil C_{\mathrm{rank}}(n-1)\rceil$ and the batch budget $T_{\max}\leftarrow\lceil C_{\mathrm{try}}n\rceil$.
7:Set a per-batch oracle-call cap $B_{\max}\leftarrow\lceil C_{\mathrm{samp}}\cdot N\rceil$.
8:Precompute $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$ $\triangleright$ Exists by AC2-AC3. Used inside AC4.
9:for $\mathrm{trial}=1$ to $T_{\max}$ do
10:   Initialize $\mathcal{S}\leftarrow\emptyset$ and $\mathrm{calls}\leftarrow 0$.
11:  while $|\mathcal{S}|<N$ and $\mathrm{calls}<B_{\max}$ do
12:     $\mathrm{calls}\leftarrow\mathrm{calls}+1$.
13:     $\bm{u}_{Q}\leftarrow\textsc{Step9DaggerSample}(D,M,Q,\mathrm{inv})$ $\triangleright$ Algorithm 2.
14:   if $\bm{u}_{Q}\neq\bm{0}$ then
15:        Append $\bm{u}_{Q}$ as a row of $\mathcal{S}$.
16:   end if
17:  end while
18:  if $|\mathcal{S}|<N$ then
19:     continue $\triangleright$ Sampling cap exceeded (wrong guess or atypical transcript).
20:  end if
21:   $\mathbf{b}_{Q}\leftarrow\textsc{RecoverBModQ}(\mathcal{S},Q)$ $\triangleright$ Algorithm 4.
22:  if $\mathbf{b}_{Q}=\textsc{fail}$ then
23:     continue
24:  end if
25:   $(s,e)\leftarrow\textsc{ExtractAndVerifyLWE}(\mathbf{b}_{Q},\mathbf{A},q,\beta,p_{1},\dots,p_{\kappa})$ $\triangleright$ Algorithm 3.
26:  if $(s,e)\neq\textsc{fail}$ then
27:     return $(s,e)$
28:  end if
29:end for
30:return fail
Algorithm 2 Step9DaggerSample: chirp correction and direct Fourier sampling
1:Input: Parameters $(D,M,Q)$ and $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$.
2:Oracle: One copy of $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$ (Definition 2.7).
3:Output: A reduced sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ (possibly $\bm{0}$).
4:Query the state-preparation oracle to obtain $|\widetilde{\varphi}_{7}\rangle$ on $X$.
5:Measure $\tau\leftarrow X_{1}\bmod D^{2}$ $\triangleright$ Deterministic conditioned on the transcript by AC1/AC4.
6:Compute $v_{1,Q}\leftarrow\operatorname{ctr}(\tau)\bmod Q$.
7:Apply the $\tau$-controlled diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$ as defined in AC4 (using $\mathrm{inv}$ and $v_{1,Q}$).
8:Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to $X$ and measure $\bm{u}\in\mathbb{Z}_{M}^{n}$.
9:return $\bm{u}_{Q}\leftarrow\bm{u}\bmod Q$.
Algorithm 3 ExtractAndVerifyLWE: parse and certify $(s,e)$
1:Input: $\mathbf{b}_{Q}\in\mathbb{Z}_{Q}^{n}$, the defining matrix $\mathbf{A}$ of $L_{q}^{\perp}(\mathbf{A})$, and public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
2:Output: The LWE secret $s$ and error $e$, or fail.
3:Let $m\leftarrow$ the number of rows of $\mathbf{A}$ and $n\leftarrow$ the number of columns of $\mathbf{A}$.
4:Let $\ell\leftarrow n-m-1$ (so that $\mathbf{A}$ has block form $2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$ with $\mathbf{U}\in\mathbb{Z}_{q}^{\ell\times m}$ in the chosen-secret embedding).
5:Lift $\mathbf{b}_{Q}$ coordinate-wise to $\mathbf{b}\in(-Q/2,Q/2]^{n}\cap\mathbb{Z}^{n}$.
6:if $b_{1}\neq-1$ as an integer then
7:  return fail
8:end if
9:for $i=2$ to $n$ do
10:  if $2p_{1}\nmid b_{i}$ in $\mathbb{Z}$ then
11:   return fail
12:  end if
13:end for
14:for $\eta=2$ to $\kappa$ do
15:  if $b_{\eta}\neq 2p_{1}p_{\eta}$ in $\mathbb{Z}$ then
16:   return fail
17:  end if
18:end for
19:Parse $\mathbf{b}=[-1,\,2p_{1}s^{\top},\,2p_{1}e^{\top}]^{\top}$ with $s\in\mathbb{Z}^{\ell}$ and $e\in\mathbb{Z}^{m}$, and recover $s,e$ by integer division by $2p_{1}$.
20:Check shortness (e.g. $\|s\|_{\infty},\|e\|_{\infty}\leq\beta\log n$) and any promised norm bounds for $\|\mathbf{b}\|$ from the lattice analysis.
21:if $\mathbf{A}\mathbf{b}\not\equiv\bm{0}\pmod{q}$ then
22:  return fail
23:end if
24:return $(s,e)$.
Algorithm 4 RecoverBModQ: solve for $\mathbf{b}\bmod Q$ from dual samples
1:Input: A matrix $\mathcal{S}\in(\mathbb{Z}_{Q})^{N\times n}$ whose rows are nonzero samples $\bm{u}_{Q}$.
2:Output: $\mathbf{b}_{Q}=\mathbf{b}\bmod Q$, or fail.
3:Let $U\in(\mathbb{Z}_{Q})^{N\times(n-1)}$ be columns $2..n$ of $\mathcal{S}$, and let $\bm{y}\in(\mathbb{Z}_{Q})^{N}$ be column 11 of $\mathcal{S}$.
4:Solve $U\bm{z}\equiv\bm{y}\pmod{Q}$ and test uniqueness (e.g. via Smith normal form).
5:if no solution exists or the solution is not unique then
6:  return fail
7:end if
8:Let $\bm{z}$ be the unique solution and set $\mathbf{b}_{Q}\leftarrow[-1,\bm{z}^{\top}]^{\top}\in\mathbb{Z}_{Q}^{n}$.
9:return $\mathbf{b}_{Q}$.
Remark 3.1 (Classical guessing loop for the norm).
As in chen2024quantum, the small modulus is chosen as $Q=(c{+}1)\|\mathbf{b}\|^{2}$, where $\mathbf{b}$ is the (unknown) unique shortest vector of $L_{q}^{\perp}(\mathbf{A})$. We run a classical outer loop over all polynomially many candidates $B_{\mathrm{guess}}$ for $\|\mathbf{b}\|^{2}$. For each guess we set $Q\leftarrow(c{+}1)B_{\mathrm{guess}}$, choose an odd $D$ with $\gcd(D,Q)=1$, derive $M=2D^{2}Q$ (hence $M_{2}=D^{2}Q$), and complete the remaining Karst-wave parameters as in Definition 2.1. We then execute Algorithm 1 for this guess.
Only the correct guess yields dual samples supported on the hyperplane $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ and passes the explicit verification in Algorithm 3, except with negligible probability. To keep rejection for incorrect guesses polynomial-time, Algorithm 1 uses a fixed trial budget $T_{\max}=\operatorname{poly}(n)$ and a per-trial sampling cap $B_{\max}=\operatorname{poly}(n)$, returning fail if no verified solution is found within budget.
Remark 3.2 (Instantiation for LWE).
For the $q$-ary lattice $L_{q}^{\perp}(\mathbf{A})$ in chen2024quantum, the promised unique shortest vector has the form
$\mathbf{b}=[-1,\,2p_{1}s^{\top},\,2p_{1}e^{\top}]^{\top}\in\mathbb{Z}^{n},$
where $s$ is the LWE secret and $e$ is the error vector. Because $\|\mathbf{b}\|_{\infty}\leq 2p_{1}\beta\log n<Q/2$ in the parameter regime of interest, the reduction map $\mathbf{b}\mapsto\mathbf{b}\bmod Q$ is injective on the centred cube that contains the true shortest vector. Therefore, once Algorithm 1 recovers $\mathbf{b}\bmod Q$, the canonical lift to $(-Q/2,Q/2]^{n}$ recovers $\mathbf{b}$ as an integer vector. Dividing the last n1n-1 coordinates by $2p_{1}$ (after checking divisibility) yields $s$ and $e$, and the check $\mathbf{A}\mathbf{b}\equiv 0\pmod{q}$ certifies correctness.
Fourier sampling.
We apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate register block and then measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ (Algorithm 2). The outcome distribution is analyzed in the next part.
3.1 Correctness under Additional Conditions
Lemma 3.3 (Dual sampling via direct QFT).
[…]
where $\Delta=2D^{2}\mathbf{b}$ and $M_{2}=M/2$. After applying $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers and measuring $\bm{u}\in\mathbb{Z}_{M}^{n}$, we have
$\Pr\bigl[\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}\bigr]\geq 1-\operatorname{negl}(n).$
Moreover, by Lemma 2.2, conditioned on the event $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ the reduced outcome $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane $H$, hence $\Pr[\bm{u}\bmod Q=\bm{0}]\leq Q^{-(n-1)}+\operatorname{negl}(n)$.
Proof of Lemma 3.3.
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals
$\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}\cdot\sum_{j\in\mathbb{Z}}\sum_{\bm{k}\in\mathcal{K}}\alpha_{E}(j,\bm{k})\,i^{\|\bm{k}\|^{2}}\,\omega_{M}^{\,j\langle\bm{u},\Delta\rangle+M_{2}\langle\bm{u},\bm{k}\rangle},$
so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
By AC5, conditioned on $E$ the coefficients factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$. Using $M_{2}=M/2$ we have $\omega_{M}^{M_{2}\langle\bm{u},\bm{k}\rangle}=(-1)^{\langle\bm{u},\bm{k}\rangle}$, and since $k_{1}=0$ this depends only on the parity pattern of $\bm{u}_{[2..n]}$. Moreover $M=2D^{2}Q$ and $\Delta=2D^{2}\mathbf{b}$, hence
$\omega_{M}^{\,j\langle\bm{u},\Delta\rangle}=\exp\!\left(2\pi i\,j\,\frac{t(\bm{u})}{Q}\right),$
where $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and we identify $t(\bm{u})$ with its representative in $\{0,\dots,Q-1\}$ when used inside the exponential. Therefore $S_{E}(\bm{u})$ factorizes as
$S_{E}(\bm{u})=\Bigl(\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\,t(\bm{u})/Q}\Bigr)\cdot\Bigl(\sum_{\bm{k}\in\mathcal{K}}\beta_{E}(\bm{k})\,i^{\|\bm{k}\|^{2}}\,(-1)^{\langle\bm{u},\bm{k}\rangle}\Bigr)=\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)\cdot G_{E}(\bm{u}),$
where $G_{E}(\bm{u})$ depends only on the parity pattern of $\bm{u}_{[2..n]}$. Thus
$\Pr[\bm{u}]\ \propto\ |S_{E}(\bm{u})|^{2}=|\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)|^{2}\cdot|G_{E}(\bm{u})|^{2}.$
For each residue $t\in\mathbb{Z}_{Q}$, let
$T_{t}:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\langle\mathbf{b},\bm{u}\rangle\equiv t\pmod{Q}\}$.
Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, for every fixed $\bm{u}_{[2..n]}\in\mathbb{Z}_{M}^{n-1}$ and every $t$ there are exactly $M/Q=2D^{2}$ values of $u_{1}\in\mathbb{Z}_{M}$ such that $\bm{u}\in T_{t}$. Since $G_{E}(\bm{u})$ does not depend on $u_{1}$, the total grid mass $\sum_{\bm{u}\in T_{t}}|G_{E}(\bm{u})|^{2}$ is the same for every $t$. Hence the total probability mass of $T_{t}$ is proportional to $|\widehat{\alpha}_{E}(t/Q)|^{2}$, up to negligible truncation error.
By the leakage bound (2.1) (good transcript) and this proportionality, the total off-resonant probability mass satisfies
$\sum_{t\neq 0}\Pr[T_{t}]=\frac{\sum_{t\neq 0}|\widehat{\alpha}_{E}(t/Q)|^{2}}{\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}}\leq\operatorname{negl}(n),$
so
$\Pr\bigl[\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}\bigr]\geq 1-\operatorname{negl}(n).$
Finally, conditioned on this resonance event, Lemma 2.2 gives exact uniformity of $\bm{u}\bmod Q$ on $H$, and in particular $\Pr[\bm{u}\bmod Q=\bm{0}]\leq Q^{-(n-1)}+\operatorname{negl}(n)$. ∎
Theorem 3.4 (Step $9^{\dagger}$ is correct).
Assume AC1-AC5 and run Algorithm 1 in the correct iteration of the outer norm-guessing loop. Each sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ appended to $\mathcal{S}$ by Algorithm 1 is non-zero and satisfies
$\langle\mathbf{b},\bm{u}_{Q}\rangle\equiv 0\pmod{Q}$
except with probability $\operatorname{negl}(n)$ over the choice of a good transcript and the internal randomness of the measurements. Solving the resulting linear system modulo $Q$ (using the known normalization $b_{1}=-1$) recovers $\mathbf{b}\bmod Q$ and hence $\mathbf{b}$ as an integer vector. The LWE secret and error are then extracted and verified in polynomial time with overwhelming probability.
Proof.
By Additional Condition AC5, a transcript drawn from the front end is good except with probability $\operatorname{negl}(n)$. Algorithm 1 makes only $\operatorname{poly}(n)$ oracle calls, so by a union bound we may condition on the event that all invocations are good, losing at most a negligible term in the overall success probability.
In each oracle call within Algorithm 1, the sampling primitive Step9DaggerSample (Algorithm 2) measures $\tau=X_{1}\bmod D^{2}$ and applies the diagonal gate $U_{\mathrm{corr}}^{(\tau)}$. Lemma 2.4 (which uses the canonical-gauge requirement in AC4) shows that this removes the quadratic chirp and produces the chirp-free state $|\varphi_{7}\rangle$ assumed in AC5.
Applying $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ and measuring yields $\bm{u}\in\mathbb{Z}_{M}^{n}$. Lemma 3.3 implies that $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ except with probability $\operatorname{negl}(n)$, and Lemma 2.2 implies that, conditioned on this resonance event, $\bm{u}\bmod Q$ is exactly uniform on $H=\{\bm{v}\in\mathbb{Z}_{Q}^{n}:\langle\mathbf{b},\bm{v}\rangle\equiv 0\pmod{Q}\}$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}]\leq Q^{-(n-1)}+\operatorname{negl}(n)$, so the rejection step $\bm{u}_{Q}\neq\bm{0}$ fails only negligibly often. A union bound over the $N=O(n)$ accepted samples in a batch implies that, with probability $1-\operatorname{negl}(n)$, every accepted $\bm{u}_{Q}$ satisfies $\langle\mathbf{b},\bm{u}_{Q}\rangle\equiv 0\pmod{Q}$.
Fix such a successful batch. Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, the projection $H\to\mathbb{Z}_{Q}^{n-1}$ given by $\bm{u}\mapsto\bm{u}_{[2..n]}$ is a bijection. Hence the accepted vectors $\bm{u}_{Q,[2..n]}$ are i.i.d. (up to a negligible statistical error from the rare off-resonant outputs) uniform over $\mathbb{Z}_{Q}^{n-1}\setminus\{\bm{0}\}$. Let $U\in\mathbb{Z}_{Q}^{N\times(n-1)}$ be the matrix with these rows and let $\bm{y}$ be the vector of first coordinates. Then the true unknown $\mathbf{b}_{[2..n]}\bmod Q$ satisfies $U\mathbf{b}_{[2..n]}\equiv\bm{y}\pmod{Q}$.
With overwhelming probability $U$ is injective as a map $(\mathbb{Z}_{Q})^{n-1}\to(\mathbb{Z}_{Q})^{N}$. Indeed, for any prime $p\mid Q$, the reduction $U\bmod p$ is (up to a negligible error from excluding the all-zero row) an i.i.d. uniform random matrix over the field $\mathbb{Z}_{p}$ with $N\geq C_{\mathrm{rank}}(n-1)$ rows, and standard rank bounds give $\Pr[\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)<n-1]\leq p^{-\Omega(n)}$. If $\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)=n-1$ and $U\bm{z}\equiv\bm{0}\pmod{p^{e}}$, then reducing modulo $p$ gives $\bm{z}\equiv\bm{0}\pmod{p}$; writing $\bm{z}=p\bm{z}_{1}$ and iterating shows $\bm{z}\equiv\bm{0}\pmod{p^{e}}$, so $\ker(U\bmod p^{e})=\{\bm{0}\}$. By the Chinese remainder theorem, $\ker(U\bmod Q)=\{\bm{0}\}$. Therefore the system has a unique solution $\bm{z}\in\mathbb{Z}_{Q}^{n-1}$, which Algorithm 4 computes (e.g. via Smith normal form) and uses to set $\mathbf{b}_{Q}=[-1,\bm{z}^{\top}]^{\top}=\mathbf{b}\bmod Q$.
Finally, the bound $\|\mathbf{b}\|_{\infty}<Q/2$ from (R1) makes the centered lift from $\mathbb{Z}_{Q}^{n}$ to $(-Q/2,Q/2]^{n}$ injective on the relevant region, so lifting $\mathbf{b}_{Q}$ recovers $\mathbf{b}$ as an integer vector. The subsequent divisibility, planted-prefix, shortness, and lattice-membership checks (Algorithm 3) succeed for the correct guess and reject incorrect guesses, as in chen2024quantum. ∎
Corollary 3.5 (Conditional quantum algorithm for LWE).
Assume the LWE-to-lattice reduction and the unique-shortest-vector promise for $L_{q}^{\perp}(\mathbf{A})$ from chen2024quantum. If the corresponding front end can be compiled in a re-tuned configuration satisfying Additional Conditions AC1-AC5, then Algorithm 1 solves the chosen-secret LWE instance encoded by $L_{q}^{\perp}(\mathbf{A})$ in quantum polynomial time with overwhelming probability. Via the classical reductions of chen2024quantum, this yields a quantum algorithm for standard LWE in the same parameter regime, conditional on AC1-AC5.
4 Conclusion
In this work, we present a new Step $9^{\dagger}$, which turns unknown offsets into harmless phase patterns. Under explicit front-end conditions on the Karst wave window, it yields modular linear relations from direct Fourier sampling on the coordinate registers. We instantiate this step with the parameters and the lattice $L_{q}^{\perp}(\mathbf{A})$ from chen2024quantum. In the access model of Section 2.2, such a construction gives a direct path from a superposition with support $\{j\Delta+\bm{v}^{\ast}\}$ to samples from the dual lattice. Under these assumptions, we obtain a quantum algorithm that solves the chosen-secret LWE problem of chen2024quantum in almost the same parameter regime as in that work.
Acknowledgment
We are grateful to all who provided constructive discussions and helpful feedback.

### v7

@@
 We revisit the post-processing phase of Chen’s Karst-wave quantum lattice algorithm (chen2024quantum) in the Learning with Errors (LWE) parameter regime.
-Conditioned on a transcript $E$, the post-Step 7 coordinate state on $(\mathbb{Z}_{M})^{n}$ is supported on the shifted grid line
+Conditioned on a transcript $E$, the post-Step 7 coordinate state on $(\mathbb{Z}_{M})^{n}$ is supported on an affine grid line
 $\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\},$
-where $\Delta=2D^{2}\mathbf{b}$ and $M=2M_{2}=2D^{2}Q$ with $Q$ odd.
-The amplitudes contain a quadratic chirp $\omega_{Q}^{-j^{2}}$ and an unknown run-dependent offset $\bm{v}^{\ast}(E)$.
-We show that Chen’s Steps 8-9 can be replaced by a single routine, Step $9^{\dagger}$: measure $\tau:=X_{1}\bmod D^{2}$, apply a $\tau$-dependent quadratic phase on the first coordinate register to remove the chirp, and then apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers.
-The chirp removal uses a canonical gauge promise: conditioned on $E$, the offset $v_{1}^{\ast}(E)\bmod M_{2}$ equals the centered lift $\operatorname{ctr}(\tau)$.
-Under explicit front-end conditions (AC1-AC5), a Fourier sample $\bm{u}\in\mathbb{Z}_{M}^{n}$ satisfies the resonance $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ with probability $1-\operatorname{negl}(n)$, and conditioned on resonance the reduced sample $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane.
-This yields exact recovery of $\mathbf{b}\bmod Q$ and hence the embedded LWE secret.
+with $\Delta=2D^{2}\mathbf{b}$, $M=2M_{2}=2D^{2}Q$, and $Q$ odd.
+The amplitudes include a quadratic Karst-wave chirp $\omega_{Q}^{-j^{2}}$ and an unknown run-dependent offset $\bm{v}^{\ast}(E)$.
+We show that Chen’s Steps 8-9 can be replaced by a single exact post-processing routine: measure the deterministic residue $\tau:=X_{1}\bmod D^{2}$, obtain the run-local class $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ as explicit side information in our access model, apply a $v_{1,Q}$-dependent diagonal quadratic phase on $X_{1}$ to cancel the chirp, and then apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the coordinate registers.
+The routine never needs the full offset $\bm{v}^{\ast}(E)$.
+Under Additional Conditions AC1-AC5 on the front end, a measured Fourier outcome $\bm{u}\in\mathbb{Z}_{M}^{n}$ satisfies the resonance $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$ with probability $1-\operatorname{negl}(n)$.
+Moreover, conditioned on resonance, the reduced outcome $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane $H=\{\bm{v}\in\mathbb{Z}_{Q}^{n}:\langle\mathbf{b},\bm{v}\rangle\equiv 0\pmod{Q}\}$.
 1 Introduction
@@
 […]
-for an outcome-dependent amplitude profile $\alpha_{E}(j,\bm{k})$ and an index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$ that comes from the grid-state construction.
+for an outcome-dependent amplitude profile $\alpha_{E}(j,\bm{k})$ and an index set $\mathcal{K}\subseteq\{0\}\times\{0,1\}^{n-1}$ that comes from the grid-state construction.
 When it is clear from context, we suppress the transcript index $E$.
@@
 We view the coordinate values as elements of $(\mathbb{Z}_{M})^{n}$ throughout.
-Accordingly, $\mathbf{X}(j):=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ is interpreted componentwise modulo $M$ (as in Chen’s post-Step 7 expression), and $\bm{k}$ records the $M_{2}$-shift term $M_{2}\bm{k}$.
+Accordingly, $\mathbf{X}(j):=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ is interpreted componentwise modulo $M$, and $\bm{k}$ records the $M_{2}$-shift term $M_{2}\bm{k}$.
 Reducing the coordinate registers modulo $M_{2}$ eliminates this shift and yields the affine congruence $\mathbf{X}(j)\equiv 2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}\pmod{M_{2}}$ used in Additional Condition AC1.
@@
 […]
-Additional Condition AC1 records the affine relation $\mathbf{X}(j)=2D^{2}j\,\mathbf{b}+\bm{v}^{\ast}$ for the post-Step 7 state.
-Additional Condition AC4 fixes a canonical gauge for the first-coordinate offset: the preparation circuit is compiled (by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$) so that measuring $\tau:=X_{1}\bmod D^{2}$ determines $v_{1}^{\ast}(E)\bmod Q$ for that run.
+Additional Condition AC4 isolates the only extra access needed to cancel the Karst-wave chirp using only the first coordinate register: after measuring the deterministic residue $\tau:=X_{1}\bmod D^{2}$, the post-processing must obtain the run-local class $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$.
+This information is not implied by AC1 alone; in our access model, it is provided as a short run-local gauge value.
+If the gauge value is available coherently inside the preparation, one can compile a canonical-gauge variant with $\lambda(E)=0$ (Proposition 2.3), in which case $v_{1,Q}=\operatorname{ctr}(\tau)\bmod Q$.
 Under Additional Condition AC1, the first coordinate satisfies the precise congruence
@@
 $X_{1}\bmod D^{2}\equiv v_{1}^{\ast}\bmod D^{2},\qquad(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$,
-where under the canonical gauge (AC4) the post-processing computes $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from $\tau=X_{1}\bmod D^{2}$.
-In Chen’s post-processing one ultimately outputs non-zero Fourier samples in $\mathbb{Z}_{M_{2}}^{n}$ (with $M_{2}=M/2$) whose reduction modulo $Q=M/(2D^{2})$ satisfies a homogeneous dual relation.
-In our direct-QFT variant we instead measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ and store only the reduction $\bm{u}_{Q}:=\bm{u}\bmod Q$.
+where under AC4 the post-processing computes $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from $\tau=X_{1}\bmod D^{2}$ together with the run-local gauge value.
+In Chen’s post-processing, one ultimately outputs non-zero Fourier samples in $\mathbb{Z}_{M_{2}}^{n}$ (with $M_{2}=M/2$) whose reduction modulo $Q=M/(2D^{2})$ satisfies a homogeneous dual relation.
+In our direct-QFT variant, we instead measure $\bm{u}\in\mathbb{Z}_{M}^{n}$ and store only the reduction $\bm{u}_{Q}:=\bm{u}\bmod Q$.
 In either view, the goal is to obtain non-zero vectors satisfying
@@
 where $Q:=\frac{M}{2D^{2}}$ is an odd integer.
-(In chen2024quantum the parameters are additionally arranged so that $Q$ has a convenient factorization, but Step $9^{\dagger}$ does not use it.) The short vector $\mathbf{b}=(b_{1},\dots,b_{n})\in\mathbb{Z}^{n}$ stores the hidden data and satisfies the normalization $b_{1}=-1$.
+The short vector $\mathbf{b}=(b_{1},\dots,b_{n})\in\mathbb{Z}^{n}$ stores the hidden data and satisfies the normalization $b_{1}=-1$.
 The offset $\bm{v}^{\ast}$ is unknown and can change from run to run.
@@
 $\langle\bm{a},\bm{u}\rangle\bmod Q\ :=\ \sum_{i=1}^{n}(a_{i}\bmod Q)\cdot(u_{i}\bmod Q)\ \in\ \mathbb{Z}_{Q}.$
-When we write $\langle\bm{a},\bm{u}\rangle\equiv 0\pmod{Q}$ this is the intended meaning (i.e., the dot product after coordinate-wise reduction $\mathbb{Z}_{M}\to\mathbb{Z}_{Q}$).
+When we write $\langle\bm{a},\bm{u}\rangle\equiv 0\pmod{Q}$, this is the intended meaning.
 Representative conventions.
@@
 $0\!\mid\!\mathbb{Z}^{n-1}:=\{0\}\times\mathbb{Z}^{n-1}.$
+We also use
+$0\!\mid\!\{0,1\}^{n-1}:=\{0\}\times\{0,1\}^{n-1}.$
+[…]
 The physical coordinate block in Chen’s Step 7 lives in $\mathbb{Z}_{M}$ with $M=2M_{2}$.
@@
 We use $\operatorname{poly}(\cdot)$ for an unspecified fixed polynomial and $\operatorname{negl}(n)$ for a negligible function of $n$.
+In the intended LWE instantiation (Definition 2.1) we have $Q=\operatorname{poly}(n)$, hence $Q^{-(n-1)}=\operatorname{negl}(n)$.
 Idealized unitaries vs. circuit precision.
-For clarity we treat $\mathrm{QFT}_{\mathbb{Z}_{M}}$ and the controlled root-of-unity phases used in the chirp correction (AC4) as exact unitaries.
-In a standard uniform gate set these operations can be implemented to precision $2^{-\operatorname{poly}(n)}$, which perturbs all output distributions by at most $\operatorname{negl}(n)$ in total variation distance.
+For clarity, we treat $\mathrm{QFT}_{\mathbb{Z}_{M}}$ and the controlled root-of-unity phases used in the chirp correction (AC4) as exact unitaries.
+In a standard uniform gate set, these operations can be implemented to precision $2^{-\operatorname{poly}(n)}$, which perturbs all output distributions by at most $\operatorname{negl}(n)$ in total variation distance.
 We suppress this approximation issue throughout, since it does not affect any of the modular-algebraic arguments.
+See, e.g., standard treatments of approximate QFT and fault-tolerant synthesis (nielsen2010quantum).
 Finally, we write $\operatorname{ctr}:\mathbb{Z}_{D^{2}}\to\mathbb{Z}$ for the centered lift: for $\tau\in\mathbb{Z}_{D^{2}}$, $\operatorname{ctr}(\tau)$ is the unique integer in $(-D^{2}/2,D^{2}/2]$ that is congruent to $\tau$ modulo $D^{2}$.
@@
 We follow this identification.
-Chen’s Step 9 defines a relabelled vector $\mathbf{b}^{\ast}$ (Eq.
-(39) in that paper).
+Chen’s Step 9 defines a relabelled vector $\mathbf{b}^{\ast}$.
 Our post-processing never uses $\mathbf{b}^{\ast}$.
 We keep all formulas in terms of $\mathbf{b}$ instead.
-Under Chen’s parameter constraints the quantity $M/(2D^{2})$ is an odd integer. We write
+Under Chen’s parameter constraints, the quantity $M/(2D^{2})$ is an odd integer. We write
 $Q:=\frac{M}{2D^{2}},\qquad M_{2}:=\frac{M}{2}=D^{2}Q.$
-In Chen’s concrete instantiation $Q$ happens to admit a factorisation into pairwise coprime odd factors; the new Step $9^{\dagger}$ does not use this extra structure and only relies on $Q$ being odd.
+In Chen’s concrete instantiation, $Q$ happens to admit a factorization into pairwise coprime odd factors, the new Step $9^{\dagger}$ does not use this extra structure and only relies on $Q$ being odd.
 Chen uses the letter $P$ for a large loop modulus with $P\approx M^{2}/2$. We keep this convention and reserve $P$ only for that modulus.
@@
 The only additional analytic requirement used by Step $9^{\dagger}$ is the one-dimensional spectral concentration property recorded in Additional Condition AC5, which is achieved in Chen’s complex-Gaussian regime once the loop-envelope width satisfies $\sigma_{J}\gtrsim Q\log n$ (Remark 2.5).
-All subsequent correctness statements are conditional on the existence of a (re-tuned) front-end configuration satisfying Definition 2.1 and Additional Conditions AC1-AC5.
+All subsequent correctness statements are conditional on the existence of a re-tuned front-end configuration satisfying Definition 2.1 and Additional Conditions AC1-AC5.
 Definition 2.1 (Re-tuned Karst-wave configuration).
 Take LWE parameters $(\ell,m,q,\beta)$ in the regime of chen2024quantum and set $n:=1+\ell+m$.
-Let $L_{q}^{\perp}(\mathbf{A})$, the planted shortest vector $\mathbf{b}$, and the integers $(p_{\eta})_{\eta}$ be as in Section 3.2 of chen2024quantum.
+Let $L_{q}^{\perp}(\mathbf{A})$, the planted shortest vector $\mathbf{b}$, and the integers $(p_{\eta})_{\eta}=(p_{1},\dots,p_{\kappa})$ be as in Section 3.2 of chen2024quantum.
+We assume $\kappa\geq 2$ and $\kappa-1\leq\ell$, so that the $(\kappa-1)$ chosen secret coordinates fit in the $\ell$-dimensional secret.
 A re-tuned Karst-wave configuration is any choice of front-end parameters
@@
   3. (R3)
-    Choose $r,s>0$ so that Chen’s Karst-wave constraints C.5-C.7 hold and the resulting geometric width parameter
+    Choose $r,s>0$ so that Chen’s Karst-wave constraints C.5-C.7 hold, and the resulting geometric width parameter
     $\sigma_{\mathrm{geom}}:=\frac{rs^{2}t}{u\sqrt{r^{4}+s^{4}}}$
@@
   4. (R4)
-    For the resulting complex-Gaussian front end, let $\sigma_{J}$ denote the standard deviation of the one-dimensional envelope $\alpha_{E}(j)$ along the loop index $j$ (after marginalizing over the grid index and the workspace and conditioning on a transcript $E$).
+    For the resulting complex-Gaussian front end, let $\sigma_{J}$ denote the standard deviation of the one-dimensional envelope $\alpha_{E}(j)$ along the loop index $j$.
 We restrict to parameter choices satisfying
     $\sigma_{J}\geq C_{\mathrm{spec}}\,Q\log n$
-    for some fixed absolute constant $C_{\mathrm{spec}}>0$. A lower bound of this form keeps the spectral leakage in Additional Condition AC5 bounded by $\exp(-\Omega(\log^{2}n))$.
+    for some fixed absolute constant $C_{\mathrm{spec}}>0$.
+A lower bound of this form keeps the spectral leakage in Additional Condition AC5 bounded by $\exp(-\Omega(\log^{2}n))$.
+In contrast, the geometric width parameter $\sigma_{\mathrm{geom}}$ in (R3) controls the within-ball width of each extracted Gaussian ball in coordinate space; these are distinct quantities in Chen’s regime.
 When we refer to Chen’s front end with re-tuned parameters below, we mean an outcome-conditioned circuit family compiled with a fixed choice of parameters that satisfies (R1)-(R4).
@@
 $\bigl\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\bigr\},$
-for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$.
+for some index set $\mathcal{K}\subseteq 0\!\mid\!\{0,1\}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$.
+As above, $\mathcal{K}$ is a choice of parity representatives; any larger grid index domain can be reduced to this form by regrouping all terms with the same $\bm{k}\bmod 2$ into a single amplitude.
 Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields the affine congruence
@@
 We assume throughout that $Q$ is an odd integer.
-The post-processing in Step $9^{\dagger}$ works for an arbitrary composite modulus $Q$; whenever we invert a residue class we only do so modulo a modulus that is guaranteed to be coprime to that residue (in particular, we only invert $2D^{2}$ modulo $Q$, which is valid by AC2-AC3).
+The post-processing in Step $9^{\dagger}$ works for an arbitrary composite modulus $Q$.
+Whenever we invert a residue class, we only do so modulo a modulus that is guaranteed to be coprime to that residue.
 In particular, we do not require $Q$ to factor in any prescribed way, and the integers $(p_{\eta})_{\eta}$ from the LWE-to-lattice reduction play no role in the algebraic structure of $Q$.
@@
 AC4.
-The post-Step 7 amplitudes contain the quadratic chirp $\omega_{Q}^{-j^{2}}$ in the loop index $j$. We cancel it with a diagonal unitary acting only on the first coordinate register $X_{1}$.
-Fix a transcript $E$. By AC1 and $b_{1}=-1$, for every $j$ in the effective window,
-$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}}$.
-Because $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$, reducing modulo $D^{2}$ removes the $j$-term on every branch:
+Run-local chirp cancellation from the first coordinate.
+The post-Step 7 amplitudes contain the quadratic chirp $\omega_{Q}^{-j^{2}}$ in the loop index $j$.
+Fix a transcript $E$.
+By AC1 and $b_{1}=-1$, for every $j$ in the effective window,
+$X_{1}\equiv-2D^{2}j+v_{1}^{\ast}(E)\pmod{M_{2}},\qquad M_{2}=D^{2}Q.$
+Since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$, reducing modulo $D^{2}$ removes the $j$-term on every branch:
 $\tau:=X_{1}\bmod D^{2}\equiv v_{1}^{\ast}(E)\bmod D^{2}.$
-Thus $\tau$ is deterministic conditioned on $E$; measuring it does not disturb the coherent superposition over $j$ and $\bm{k}$.
-Let $\bar{v}_{1}:=\operatorname{ctr}(\tau)\in(-D^{2}/2,D^{2}/2]$ and set $v_{1,Q}:=\bar{v}_{1}\bmod Q\in\mathbb{Z}_{Q}$.
-Under the canonical-gauge requirement below, this $v_{1,Q}$ equals $v_{1}^{\ast}(E)\bmod Q$ for that run.
-Canonical gauge (required).
-Measuring $\tau$ determines $v_{1}^{\ast}(E)\bmod D^{2}$ but leaves an ambiguity by multiples of $D^{2}$ inside $\mathbb{Z}_{M_{2}}$.
-Equivalently, there is a unique $\lambda(E)\in\mathbb{Z}_{Q}$ such that
-$v_{1}^{\ast}(E)\equiv\bar{v}_{1}+\lambda(E)D^{2}\pmod{M_{2}}$.
-Any correction depending only on $\tau$ cancels $\omega_{Q}^{-j^{2}}$ only up to a residual run-dependent linear phase $\omega_{Q}^{-\lambda(E)j}$ (up to a global constant), which shifts the resonance class $\langle\mathbf{b},\bm{u}\rangle\bmod Q$ and prevents aggregating samples across invocations.
-We therefore assume a canonical gauge in which $\lambda(E)=0$ for every transcript $E$, i.e.,
-$v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$.
-When the preparation circuit is known, this can be enforced by an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$, provided $E$ allows reversible computation of $\lambda(E)$; see Proposition 2.3.
-Define the classical inverse $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$, which exists by AC2-AC3. For each basis value $x\in\mathbb{Z}_{M}$ of $X_{1}$, write $x_{Q}:=x\bmod Q$ and define
-$j_{x}:=\mathrm{inv}\cdot\bigl((x_{Q}-v_{1,Q})\bmod Q\bigr)\in\mathbb{Z}_{Q}.$
-For a branch with loop label $j$ we have $(x_{Q}-v_{1,Q})\equiv-2D^{2}j\pmod{Q}$, hence $j_{x}\equiv j\pmod{Q}$. We apply the diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$:
+Thus $\tau$ is deterministic conditioned on $E$; measuring it does not disturb the coherent superposition over $j$ and $\bm{k}$. Let $\bar{v}_{1}:=\operatorname{ctr}(\tau)\in(-D^{2}/2,D^{2}/2]$.
+Side information. We assume the preparation oracle also returns, as classical side information, either the residue
+$v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q\in\mathbb{Z}_{Q},$
+or equivalently a “high-order” gauge value $\lambda(E)\in\mathbb{Z}_{Q}$ satisfying $v_{1}^{\ast}(E)\equiv\bar{v}_{1}+\lambda(E)D^{2}\pmod{M_{2}}$, so that $v_{1,Q}=(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$.
+(See Proposition 2.3 for a canonical-gauge compilation where $\lambda(E)=0$.)
+Chirp-cancellation unitary. Define $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$, which exists by AC2-AC3. For each basis value $x\in\mathbb{Z}_{M}$ of $X_{1}$, write $x_{Q}:=x\bmod Q$ and define
+[…]
+We apply the diagonal unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on $X_{1}$,
 […]
@@
 We extend $\alpha_{E}(j,\bm{k})$ by zero outside an interval $\{|j|\leq J_{\max}\}$.
-The effective window satisfies $J_{\max}=\operatorname{poly}(Q\log n)\ll P$ in the Karst-wave regime below; in particular we assume $J_{\max}<P/2$ so that the embedding of $j\in\mathbb{Z}_{P}$ into integers is unambiguous on the support of the state.
+The effective window satisfies $J_{\max}=\operatorname{poly}(Q\log n)\ll P$ in the Karst-wave regime below; in particular, we assume $J_{\max}<P/2$ so that the embedding of $j\in\mathbb{Z}_{P}$ into integers is unambiguous on the support of the state.
 The infinite sum above then matches, up to negligible error, the finite sum that appears in Chen’s state.
-As in Chen’s Karst-wave front end, we assume that conditioned on $E$ the amplitudes factor across the loop index and the grid label:
+As in Chen’s Karst-wave front end, we assume that conditioned on $E$, the amplitudes factor across the loop index and the grid label:
 $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\,\beta_{E}(\bm{k}).$
@@
 Additional Condition AC1 is the standard offset-coherence invariant of Chen’s deferred-measurement front end (Lemma 2.8).
-Additional Condition AC4 is the only genuinely new access-model promise used by Step $9^{\dagger}$: it replaces Chen’s Step 8 offset-extraction gadget by a gauge-fixing guarantee (or coherent access to the gauge parameter) sufficient to cancel the quadratic chirp using only $X_{1}$.
-It is enforceable by a compile-time basis translation on the first coordinate whenever the preparation circuit supports reversible computation of the gauge parameter $\lambda(E)$ (Proposition 2.3).
+Additional Condition AC4 is the only genuinely new access-model requirement used by Step $9^{\dagger}$: it asks for run-local access to $v_{1,Q}(E)$ (equivalently, to a gauge value $\lambda(E)$) so that the quadratic chirp can be cancelled using only $X_{1}$.
+This information is not implied by AC1 and is treated as explicit side information returned by the oracle; if the gauge value is available coherently inside the preparation, one can also compile a canonical gauge via an $E$-controlled basis translation on $X_{1}$ (Proposition 2.3).
 Finally, Additional Condition AC5 follows from the complex discrete-Gaussian Karst-wave envelope once $\sigma_{J}\geq C_{\mathrm{spec}}Q\log n$ (Proposition 2.6).
@@
 $R:=\{\bm{u}\in\mathbb{Z}_{M}^{n}:\langle\mathbf{b},\bm{u}\rangle\equiv 0\!\!\!\pmod{Q}\},\qquad H:=\{\bm{v}\in\mathbb{Z}_{Q}^{n}:\langle\mathbf{b},\bm{v}\rangle\equiv 0\!\!\!\pmod{Q}\}$.
-Conditioned on $R$, the reduced outcome $\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ is exactly uniform on $H$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=1/|H|=Q^{-(n-1)}$.
+If $\Pr_{\mu_{E}}[R]>0$, then conditioned on $R$ the reduced outcome $\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ is exactly uniform on $H$. In particular, $\Pr[\bm{u}\bmod Q=\bm{0}\mid R]=1/|H|=Q^{-(n-1)}$.
 Proof.
-Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals
+Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the unnormalized amplitude equals
 $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}\cdot\sum_{j\in\mathbb{Z}}\sum_{\bm{k}\in\mathcal{K}}\alpha_{E}(j,\bm{k})\,i^{\|\bm{k}\|^{2}}\,\omega_{M}^{\,j\langle\bm{u},\Delta\rangle+M_{2}\langle\bm{u},\bm{k}\rangle},$
-so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
+so the outcome probabilities are proportional to the squared magnitude of the inner sum.
 Because $M_{2}=M/2$, we have $\omega_{M}^{M_{2}\langle\bm{u},\bm{k}\rangle}=(-1)^{\langle\bm{u},\bm{k}\rangle}$, which depends only on the parity pattern of $\bm{u}_{[2..n]}$ since $k_{1}=0$ for all $\bm{k}\in\mathcal{K}$.
-On the event $R$ we have $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, hence
-$\langle\bm{u},\Delta\rangle=2D^{2}\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{2D^{2}Q}=0\pmod{M}$,
-and therefore $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=1$ for all $j$.
+Let $\bm{u}_{Q}:=\bm{u}\bmod Q\in\mathbb{Z}_{Q}^{n}$ and set $t(\bm{u}):=\langle\mathbf{b},\bm{u}_{Q}\rangle\in\mathbb{Z}_{Q}$.
+We can reduce $\bm{u}$ modulo $Q$ inside the phase because of the special form of $\Delta$.
+Write $\bm{u}=\bm{u}_{Q}+Q\bm{t}$ with $\bm{t}\in\mathbb{Z}^{n}$ coordinatewise.
+Since $\Delta=2D^{2}\mathbf{b}$ and $M=2D^{2}Q$, we have
+$\langle Q\bm{t},\Delta\rangle=Q\cdot 2D^{2}\langle\bm{t},\mathbf{b}\rangle=M\cdot\langle\bm{t},\mathbf{b}\rangle\in M\mathbb{Z},$
+so $\omega_{M}^{\,j\langle Q\bm{t},\Delta\rangle}=1$ for all integers $j$. Therefore
+$\omega_{M}^{\,j\langle\bm{u},\Delta\rangle}=\omega_{M}^{\,j\langle\bm{u}_{Q},\Delta\rangle}=\exp\!\left(2\pi i\,j\,\frac{t(\bm{u})}{Q}\right).$
+On the event $R$ we have $t(\bm{u})=0$, hence $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=1$ for all $j$.
 Therefore, conditioned on $R$, the outcome probabilities depend on $\bm{u}$ only through the parity pattern of $\bm{u}_{[2..n]}$.
@@
 Since $Q$ is odd, as $\bm{t}$ ranges over $\mathbb{Z}_{2D^{2}}^{n}$ the parity pattern of $(\bm{v}+Q\bm{t})_{[2..n]}$ is exactly uniform over $\{0,1\}^{n-1}$, independently of $\bm{v}$.
-Hence each parity pattern occurs in $F_{\bm{v}}$ with the same multiplicity for every $\bm{v}\in H$, and since the conditional probabilities depend only on this parity pattern, $\mu_{E}[F_{\bm{v}}\mid R]$ is constant over $\bm{v}\in H$.
+Hence, each parity pattern occurs in $F_{\bm{v}}$ with the same multiplicity for every $\bm{v}\in H$, and since the conditional probabilities depend only on this parity pattern, $\mu_{E}[F_{\bm{v}}\mid R]$ is constant over $\bm{v}\in H$.
 This proves uniformity of $\bm{u}\bmod Q$ on $H$ conditioned on $R$.
@@
 ∎
-Proposition 2.3 (Canonical-gauge compilation given coherent transcript access).
-Assume AC1 and suppose the state-preparation procedure is given as an explicit circuit that admits a deferred-measurement implementation producing a coherent transcript register $E$ from which $v_{1}^{\ast}(E)\bmod M_{2}$ (equivalently, the gauge parameter $\lambda(E)$ defined in AC4) can be computed reversibly.
-Then one can compile the preparation by appending an $E$-controlled basis translation on $X_{1}$ by a multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, the resulting offset satisfies the canonical-gauge requirement $v_{1}^{\ast}(E)\equiv\operatorname{ctr}(X_{1}\bmod D^{2})\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$).
-If the preparation is a black-box state oracle that does not expose such coherent access to $\lambda(E)$, the canonical-gauge condition is an additional oracle promise and is not implied by AC1 alone.
+Proposition 2.3 (Gauge side information and canonical-gauge compilation).
+Assume AC1 and suppose the state-preparation procedure is given as an explicit circuit (or oracle) that outputs the coordinate registers together with a classical side-information register $E$ from which the gauge parameter $\lambda(E)\in\mathbb{Z}_{Q}$ of AC4 (equivalently, $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$) can be obtained.
+Then Additional Condition AC4 holds: the post-processing can compute $v_{1,Q}$ for the current run and implement the chirp-cancellation unitary on $X_{1}$.
+Moreover, if $E$ is available coherently during the preparation, one can compile the preparation by appending an $E$-controlled basis translation on $X_{1}$ by $\lambda(E)D^{2}$ (modulo $M$) so that the resulting state satisfies the canonical gauge $\lambda(E)=0$.
+If the preparation is a black-box state oracle that outputs only the coordinate registers and does not expose $\lambda(E)$ (or $v_{1}^{\ast}(E)\bmod Q$) as side information, then AC4 becomes an additional oracle promise and is not implied by AC1 alone.
 Proof.
@@
 Assume AC1-AC4.
-Then the procedure in AC4 defines, for each measured $\tau$, a unitary $U_{\mathrm{corr}}^{(\tau)}$ on the first coordinate register such that $|\varphi_{7}\rangle:=U_{\mathrm{corr}}|\widetilde{\varphi}_{7}\rangle$ has the form (1.2).
+Then the procedure in AC4 defines, for each run-local residue $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ (obtainable from $\tau$ and/or transcript side information), a unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on the first coordinate register such that $|\varphi_{7}\rangle:=U_{\mathrm{corr}}^{(v_{1,Q})}|\widetilde{\varphi}_{7}\rangle$ has the form Equation (1.2).
 The quadratic chirp $e^{-2\pi ij^{2}/Q}$ in the loop index is removed on every branch.
@@
 Conditioned on the transcript $E$, the measurement of $\tau:=X_{1}\bmod D^{2}$ is deterministic and yields $\tau\equiv v_{1}^{\ast}\bmod D^{2}$ on every branch.
-By the gauge choice in AC4, with $\bar{v}_{1}=\operatorname{ctr}(\tau)$ and $v_{1,Q}=\bar{v}_{1}\bmod Q$, we have $(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$ on every branch.
-The reversible computation of $j_{x}$ in AC4 is well-defined modulo $Q$ by AC2-AC3.
-For $x=X_{1}$ on a branch with loop label $j$ we have $j_{x}\equiv j\pmod{Q}$.
+By AC4, the post-processing obtains $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$, hence $(X_{1}\bmod Q)-v_{1,Q}\equiv-2D^{2}j\pmod{Q}$ on every branch.
+The reversible computation of $j(X_{1})$ in AC4 is well-defined modulo $Q$ by AC2-AC3.
+For $x=X_{1}$ on a branch with loop label $j$ we have $j(X_{1})\equiv j\pmod{Q}$.
 The unitary $U_{\mathrm{corr}}$ multiplies that branch by $\exp(2\pi ij^{2}/Q)$.
@@
 The map is diagonal in the computational basis of $X_{1}$.
-Hence it is unitary.
+Hence, it is unitary.
 ∎
 Remark 2.5 (Justification of Additional Condition AC5).
-Additional Condition AC5 isolates the only additional front-end property used by Step $9^{\dagger}$: spectral concentration of the one-dimensional Karst-wave envelope at the resonant frequencies $t/Q$ with $t\equiv 0\pmod{Q}$.
+Additional Condition AC5 isolates the only additional front-end property used by Step $9^{\dagger}$: spectral concentration of the one-dimensional Karst-wave envelope on the $Q$-point grid, with dominant mass at the single resonant frequency $t=0$.
 In Chen’s Karst-wave front end, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ (see Eq.
@@
 and identify $t(\bm{u})$ with its standard representative in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$ on $[0,1)$.
-Taking the envelope width $\sigma_{J}\gtrsim Q\log n$ yields $|\widehat{\alpha}_{E}(t/Q)|\leq\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|$ for all integers $t\not\equiv 0\pmod{Q}$ by standard Poisson-summation estimates.
-Since $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$, this implies the leakage bound (2.1).
-Remark on linear phases.
+Taking the envelope width $\sigma_{J}\gtrsim Q\log n$ yields $|\widehat{\alpha}_{E}(t/Q)|\leq\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|$ for all integers $t\not\equiv 0\pmod{Q}$ by standard Poisson-summation / discrete-Gaussian Fourier-decay estimates (micciancio2007worst; peikert2010efficient; regev2010learning).
+Since $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$, this implies the leakage bound Equation (2.1).
 A time shift $j\mapsto j-j_{0}$ multiplies $\widehat{\alpha}_{E}(\theta)$ by a unit-modulus factor and does not affect its magnitude.
 In contrast, a nontrivial linear phase $\alpha_{E}(j)\mapsto\alpha_{E}(j)e^{2\pi i\nu j}$ shifts the spectrum $\widehat{\alpha}_{E}(\theta)\mapsto\widehat{\alpha}_{E}(\theta+\nu)$ and can move the dominant residue class away from $t=0$ on the grid $\{t/Q\}$.
-In our setting, such a run-dependent shift is exactly the residual $\omega_{Q}^{-\lambda(E)j}$ phase arising from a gauge mismatch in AC4; the canonical-gauge promise rules it out.
+In our setting, if the post-processing were to ignore the gauge term $\lambda(E)D^{2}$ and use $v_{1,Q}=\bar{v}_{1}\bmod Q$ in place of the correct value $v_{1,Q}=(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$, then the chirp-cancellation unitary would leave a residual linear phase $\omega_{Q}^{-\lambda(E)j}$ (up to a global constant) on the loop index $j$.
+This would shift the dominant residue class away from $t=0$ on the $Q$-point grid.
+Additional Condition AC4 (or canonical-gauge compilation) rules out this mismatch by requiring the exact run-local residue $v_{1,Q}$.
 No additional near-uniformity assumption is needed for the reduced samples: Lemma 2.2 shows that conditioned on $\langle\mathbf{b},\bm{u}\rangle\equiv 0\pmod{Q}$, the reduction $\bm{u}\bmod Q$ is exactly uniform on the dual hyperplane $H$, and $\Pr[\bm{u}\bmod Q=\bm{0}\mid\langle\mathbf{b},\bm{u}\rangle\equiv 0]=1/|H|=Q^{-(n-1)}$.
 Proposition 2.6 (Spectral concentration for the re-tuned Karst wave).
-For the complex-Gaussian Karst-wave front end of chen2024quantum (where, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$), instantiated with any re-tuned configuration that satisfies Definition 2.1, Additional Condition AC5 holds (i.e., the leakage bound (2.1) fails only with probability $\operatorname{negl}(n)$ over the front-end transcript).
+For the complex-Gaussian Karst-wave front end of chen2024quantum (where, conditioned on a transcript $E$, the post-Step 7 amplitudes factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$), instantiated with any re-tuned configuration that satisfies Definition 2.1, Additional Condition AC5 holds.
 Proof.
-Under Chen’s Karst-wave analysis, conditioned on a transcript $E$ the loop envelope $\alpha_{E}(j)$ is (up to a global phase, an integer shift in $j$, and negligible truncation) a discrete Gaussian of width $\sigma_{J}$, and the factorization $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ holds.
+Under Chen’s Karst-wave analysis, conditioned on a transcript $E$ the loop envelope $\alpha_{E}(j)$ is a discrete Gaussian of width $\sigma_{J}$, and the factorization $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$ holds.
 An integer shift in $j$ does not affect $|\widehat{\alpha}_{E}(\cdot)|$.
-Any additional run-dependent linear phase on the $1/Q$ grid would shift the dominant residue class away from $t=0$; in our setting this is precisely the gauge-mismatch phenomenon discussed and ruled out in AC4.
-Thus it suffices to bound the centered discrete-Gaussian prototype $g_{\sigma}(j):=\exp(-\pi j^{2}/\sigma^{2}).$ Its Fourier series satisfies the Poisson-summation identity
+Any additional run-dependent linear phase on the $1/Q$ grid would shift the dominant residue class away from $t=0$; in our setting, this corresponds to using an incorrect $v_{1,Q}$ in the chirp correction, which AC4 rules out by supplying the run-local gauge value.
+Thus, it suffices to bound the centered discrete-Gaussian prototype $g_{\sigma}(j):=\exp(-\pi j^{2}/\sigma^{2}).$ Its Fourier series satisfies the Poisson-summation identity (see, e.g., (micciancio2007worst))
 $\sum_{j\in\mathbb{Z}}g_{\sigma}(j)\,e^{2\pi ij\theta}=\sigma\sum_{m\in\mathbb{Z}}\exp\bigl(-\pi\sigma^{2}(\theta-m)^{2}\bigr).$
@@
 $\sum_{t\in\mathbb{Z}_{Q}\setminus\{0\}}|\widehat{\alpha}_{E}(t/Q)|^{2}\leq(Q-1)\exp(-\Omega(\log^{2}n))\cdot|\widehat{\alpha}_{E}(0)|^{2}.$
-Using $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$ proves (2.1). ∎
+Using $\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}\geq|\widehat{\alpha}_{E}(0)|^{2}$ proves Equation (2.1). ∎
 2.2 Access model
 Definition 2.7 (Access model).
-We assume (circuit-level) quantum access to an outcome-conditioned state-preparation procedure that, on each invocation, outputs one copy of the post-Step 7 state $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$.
-We additionally assume a canonical-gauge implementation as in Additional Condition AC4: inside the preparation circuit the first coordinate is translated by an $E$-dependent multiple of $D^{2}$ (modulo $M$) so that, conditioned on $E$, writing $\tau:=X_{1}\bmod D^{2}$ and $\bar{v}_{1}:=\operatorname{ctr}(\tau)$, the offset satisfies $v_{1}^{\ast}(E)\equiv\bar{v}_{1}\pmod{M_{2}}$ (equivalently, $\lambda(E)=0$ in Additional Condition AC4).
-Consequently, measuring $\tau$ allows the post-processing to compute $v_{1,Q}=\bar{v}_{1}\bmod Q$ and the chirp-correction phase for that run.
-For each invocation there exists an unobserved classical transcript $E$ such that, conditioned on $E$, the reduced state on $X$ satisfies Additional Condition AC1 with offset $\bm{v}^{\ast}(E)$ and amplitudes $\alpha_{E}(j,\bm{k})$.
-Across invocations the transcript is re-sampled, so the offset $\bm{v}^{\ast}(E)$ may change from run to run and cannot be forced to repeat.
+We assume circuit-level quantum access to an outcome-conditioned state-preparation procedure that, on each invocation, outputs one copy of the post-Step 7 state $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$, together with a classical side-information register $E$ that may include Chen’s measured strings and, for AC4, a short run-local gauge value sufficient to obtain the residue $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (equivalently, $\lambda(E)\in\mathbb{Z}_{Q}$).
+We treat $E$ as a classical output register (i.e., returned in the computational basis).
+Equivalently, the oracle may measure $E$ before returning it.
+This ensures that reading $E$ and any deterministic function of $E$ is compatible with maintaining coherence over the loop index $j$ and grid label $\bm{k}$ in $|\widetilde{\varphi}_{7}\rangle$.
+We do not assume this gauge value is derivable from the Step 7 coordinate registers alone; it is treated as explicit side information returned by the oracle, and since it is fixed conditioned on $E$ it can be measured without disturbing the superposition over $j$ and $\bm{k}$.
+As a special case, if the known preparation circuit is compiled into the canonical gauge $\lambda(E)=0$ (Proposition 2.3), then $v_{1}^{\ast}(E)\bmod Q$ can be recovered from $\tau=X_{1}\bmod D^{2}$ alone and no explicit transcript access is needed for chirp cancellation.
+For each invocation there exists a transcript $E$ such that, conditioned on $E$, the reduced state on $X$ satisfies Additional Condition AC1 with offset $\bm{v}^{\ast}(E)$ and amplitudes $\alpha_{E}(j,\bm{k})$.
+Across invocations, the transcript is re-sampled, so the offset $\bm{v}^{\ast}(E)$ may change from run to run and cannot be forced to repeat.
 The post-processing is allowed to apply any quantum circuit of size $\operatorname{poly}(n)$ to the output registers, including arithmetic in $\mathbb{Z}_{M}$ and $\mathbb{Z}_{Q}$, intermediate measurements, and $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$.
@@
 $\bigl\{\,j\Delta+\bm{v}^{\ast}(E)+M_{2}\bm{k}\bmod M:j\in\mathbb{Z},\ \bm{k}\in\mathcal{K}\,\bigr\},$
-for some index set $\mathcal{K}\subseteq\{0\}\times\mathbb{Z}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$.
+for some index set $\mathcal{K}\subseteq 0\!\mid\!\{0,1\}^{n-1}$, where $\Delta=2D^{2}\mathbf{b}$ and the offset $\bm{v}^{\ast}(E)\in(\mathbb{Z}_{M})^{n}$ depends on $E$ but not on $j$.
 Reducing modulo $M_{2}$ eliminates the $M_{2}\bm{k}$ term and yields
@@
 Proof.
-This is immediate from Chen’s post-Step 7 computational-basis form (see the excerpted Eq.
-(14)): the basis values are $2Dj\,\bm{x}+\bm{v}^{\prime}+\tfrac{M}{2}\bm{k}\bmod M$ with $\bm{k}\in 0\!\mid\!\mathbb{Z}^{n-1}$ and $\bm{x}=D\mathbf{b}$.
+This is immediate from Chen’s post-Step 7 computational-basis form: the basis values are $2Dj\,\bm{x}+\bm{v}^{\prime}+\tfrac{M}{2}\bm{k}\bmod M$ with $\bm{k}\in 0\!\mid\!\mathbb{Z}^{n-1}$ and $\bm{x}=D\mathbf{b}$.
+Since $M_{2}=M/2$, only $\bm{k}\bmod 2$ affects $\tfrac{M}{2}\bm{k}\bmod M$ (and the phase $i^{\|\bm{k}\|^{2}}$), so one may fix $\bm{k}\in 0\!\mid\!\{0,1\}^{n-1}$ as representatives.
 Substituting $\bm{x}=D\mathbf{b}$ gives $2Dj\,\bm{x}=2D^{2}j\,\mathbf{b}=j\Delta$ with $\Delta=2D^{2}\mathbf{b}$, and writing $\bm{v}^{\ast}(E):=\bm{v}^{\prime}\bmod M$ yields the claimed support in $(\mathbb{Z}_{M})^{n}$.
@@
 ∎
+Algorithm 1 Step9DaggerSample: chirp correction and direct Fourier sampling
+1:Input: Parameters $(D,M,Q)$ and $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$.
+2:Oracle: One copy of $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$ and side information sufficient to obtain $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (Definition 2.7).
+3:Output: A reduced sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ (possibly $\bm{0}$).
+4:Query the state-preparation oracle to obtain $|\widetilde{\varphi}_{7}\rangle$ on $X$ and run-local side information $E$ (including the gauge value needed for AC4).
+5:Measure $\tau\leftarrow X_{1}\bmod D^{2}$ and set $\bar{v}_{1}\leftarrow\operatorname{ctr}(\tau)$ $\triangleright$ Deterministic conditioned on $E$ (AC1); needed only if $E$ encodes $\lambda(E)$.
+6:Using $(E,\bar{v}_{1})$, compute the run-local residue $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ as in AC4:
+7:   either read $v_{1,Q}$ directly from $E$, or read $\lambda(E)\in\mathbb{Z}_{Q}$ from $E$ and set $v_{1,Q}\leftarrow(\bar{v}_{1}+\lambda(E)D^{2})\bmod Q$.
+8:Apply the diagonal unitary $U_{\mathrm{corr}}^{(v_{1,Q})}$ on $X_{1}$ as defined in AC4 (using $\mathrm{inv}$).
+9:Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to $X$ and measure $\bm{u}\in\mathbb{Z}_{M}^{n}$.
+10:return $\bm{u}_{Q}\leftarrow\bm{u}\bmod Q$.
 3 The new Step $9^{\dagger}$
 Algorithm 2 Step $9^{\dagger}$
-1:Input: A lattice instance $L_{q}^{\perp}(\mathbf{A})$ arising from a chosen-secret LWE instance as in chen2024quantum (i.e., $\mathbf{A}=[2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$) together with public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
+1:Input: A chosen-secret LWE instance $(\mathbf{U},\mathbf{t})$ as in chen2024quantum over modulus $q$ (equivalently, the lattice $L_{q}^{\perp}(\mathbf{A})$ with $\mathbf{A}=[2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$), together with public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
 2:Oracle: State-preparation access to $|\widetilde{\varphi}_{7}\rangle$ as in Definition 2.7.
-3:Parameters: Re-tuned Karst-wave configuration $(c,D,r,s,t,u,P,M,Q)$ (see Remark 3.1).
+3:Parameters: A re-tuned Karst-wave configuration $(c,D,r,s,t,u,P,M,Q)$ as in Definition 2.1 (chosen for the current norm guess; see Remark 3.1).
 4:Output: Either the LWE secret $s$ and error $e$, or fail (meaning the current norm guess is rejected).
-5:Fix absolute constants $C_{\mathrm{rank}}>1$, $C_{\mathrm{try}}\geq 1$, and $C_{\mathrm{samp}}\geq 1$.
-6:Set the target sample count $N\leftarrow\lceil C_{\mathrm{rank}}(n-1)\rceil$ and the batch budget $T_{\max}\leftarrow\lceil C_{\mathrm{try}}n\rceil$.
-7:Set a per-batch oracle-call cap $B_{\max}\leftarrow\lceil C_{\mathrm{samp}}\cdot N\rceil$.
-8:Precompute $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$ $\triangleright$ Exists by AC2-AC3. Used inside AC4.
-9:for $\mathrm{trial}=1$ to $T_{\max}$ do
-10:   Initialize $\mathcal{S}\leftarrow\emptyset$ and $\mathrm{calls}\leftarrow 0$.
-11:  while $|\mathcal{S}|<N$ and $\mathrm{calls}<B_{\max}$ do
-12:     $\mathrm{calls}\leftarrow\mathrm{calls}+1$.
-13:     $\bm{u}_{Q}\leftarrow\textsc{Step9DaggerSample}(D,M,Q,\mathrm{inv})$ $\triangleright$ Algorithm 2.
-14:   if $\bm{u}_{Q}\neq\bm{0}$ then
-15:        Append $\bm{u}_{Q}$ as a row of $\mathcal{S}$.
-16:   end if
-17:  end while
-18:  if $|\mathcal{S}|<N$ then
-19:     continue $\triangleright$ Sampling cap exceeded (wrong guess or atypical transcript).
-20:  end if
-21:   $\mathbf{b}_{Q}\leftarrow\textsc{RecoverBModQ}(\mathcal{S},Q)$ $\triangleright$ Algorithm 4.
-22:  if $\mathbf{b}_{Q}=\textsc{fail}$ then
-23:     continue
-24:  end if
-25:   $(s,e)\leftarrow\textsc{ExtractAndVerifyLWE}(\mathbf{b}_{Q},\mathbf{A},q,\beta,p_{1},\dots,p_{\kappa})$ $\triangleright$ Algorithm 3.
-26:  if $(s,e)\neq\textsc{fail}$ then
-27:     return $(s,e)$
-28:  end if
-29:end for
-30:return fail
-Algorithm 2 Step9DaggerSample: chirp correction and direct Fourier sampling
-1:Input: Parameters $(D,M,Q)$ and $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$.
-2:Oracle: One copy of $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$ (Definition 2.7).
-3:Output: A reduced sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ (possibly $\bm{0}$).
-4:Query the state-preparation oracle to obtain $|\widetilde{\varphi}_{7}\rangle$ on $X$.
-5:Measure $\tau\leftarrow X_{1}\bmod D^{2}$ $\triangleright$ Deterministic conditioned on the transcript by AC1/AC4.
-6:Compute $v_{1,Q}\leftarrow\operatorname{ctr}(\tau)\bmod Q$.
-7:Apply the $\tau$-controlled diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$ as defined in AC4 (using $\mathrm{inv}$ and $v_{1,Q}$).
-8:Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to $X$ and measure $\bm{u}\in\mathbb{Z}_{M}^{n}$.
-9:return $\bm{u}_{Q}\leftarrow\bm{u}\bmod Q$.
+5:Let $\ell\leftarrow$ the number of rows of $\mathbf{U}$ and $m\leftarrow$ the number of columns of $\mathbf{U}$; set $n\leftarrow 1+\ell+m$.
+6:Fix absolute constants $C_{\mathrm{rank}}>1$, $C_{\mathrm{try}}\geq 1$, and $C_{\mathrm{samp}}\geq 1$.
+7:Set the target sample count $N\leftarrow\lceil C_{\mathrm{rank}}(n-1)\rceil$ and the batch budget $T_{\max}\leftarrow\lceil C_{\mathrm{try}}n\rceil$.
+8:Set a per-batch oracle-call cap $B_{\max}\leftarrow\lceil C_{\mathrm{samp}}\cdot N\rceil$.
+9:Precompute $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$ $\triangleright$ Exists by AC2-AC3. Used inside AC4.
+10:for $\mathrm{trial}=1$ to $T_{\max}$ do
+11:   Initialize $\mathcal{S}\leftarrow\emptyset$ and $\mathrm{calls}\leftarrow 0$.
+12:  while $|\mathcal{S}|<N$ and $\mathrm{calls}<B_{\max}$ do
+13:     $\mathrm{calls}\leftarrow\mathrm{calls}+1$.
+14:     $\bm{u}_{Q}\leftarrow\textsc{Step9DaggerSample}(D,M,Q,\mathrm{inv})$ $\triangleright$ Algorithm 1.
+15:   if $\bm{u}_{Q}\neq\bm{0}$ then
+16:        Append $\bm{u}_{Q}$ as a row of $\mathcal{S}$.
+17:   end if
+18:  end while
+19:  if $|\mathcal{S}|<N$ then
+20:     continue $\triangleright$ Sampling cap exceeded (wrong guess or atypical transcript).
+21:  end if
+22:   $\mathbf{b}_{Q}\leftarrow\textsc{RecoverBModQ}(\mathcal{S},Q)$ $\triangleright$ Algorithm 4.
+23:  if $\mathbf{b}_{Q}=\textsc{fail}$ then
+24:     continue
+25:  end if
+26:   $(s,e)\leftarrow\textsc{ExtractAndVerifyLWE}(\mathbf{b}_{Q},\mathbf{U},\mathbf{t},q,\beta,p_{1},\dots,p_{\kappa})$ $\triangleright$ Algorithm 3.
+27:  if $(s,e)\neq\textsc{fail}$ then
+28:     return $(s,e)$
+29:  end if
+30:end for
+31:return fail
 Algorithm 3 ExtractAndVerifyLWE: parse and certify $(s,e)$
-1:Input: $\mathbf{b}_{Q}\in\mathbb{Z}_{Q}^{n}$, the defining matrix $\mathbf{A}$ of $L_{q}^{\perp}(\mathbf{A})$, and public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
+1:Input: $\mathbf{b}_{Q}\in\mathbb{Z}_{Q}^{n}$, a chosen-secret LWE instance $(\mathbf{U},\mathbf{t})$ with $\mathbf{U}\in\mathbb{Z}_{q}^{\ell\times m}$ and $\mathbf{t}\in\mathbb{Z}_{q}^{m}$, and public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
 2:Output: The LWE secret $s$ and error $e$, or fail.
-3:Let $m\leftarrow$ the number of rows of $\mathbf{A}$ and $n\leftarrow$ the number of columns of $\mathbf{A}$.
-4:Let $\ell\leftarrow n-m-1$ (so that $\mathbf{A}$ has block form $2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$ with $\mathbf{U}\in\mathbb{Z}_{q}^{\ell\times m}$ in the chosen-secret embedding).
+3:Let $\ell\leftarrow$ the number of rows of $\mathbf{U}$ and $m\leftarrow$ the number of columns of $\mathbf{U}$; set $n\leftarrow 1+\ell+m$.
+4:(Fail if $\kappa-1>\ell$.) $\triangleright$ Chosen-secret prefix must fit in the $\ell$-dimensional secret.
 5:Lift $\mathbf{b}_{Q}$ coordinate-wise to $\mathbf{b}\in(-Q/2,Q/2]^{n}\cap\mathbb{Z}^{n}$.
@@
 13:end for
-14:for $\eta=2$ to $\kappa$ do
-15:  if $b_{\eta}\neq 2p_{1}p_{\eta}$ in $\mathbb{Z}$ then
+14:for $i=2$ to $\kappa$ do
+15:  if $b_{i}\neq 2p_{1}p_{i}$ in $\mathbb{Z}$ then
 16:   return fail
@@
 20:Check shortness (e.g. $\|s\|_{\infty},\|e\|_{\infty}\leq\beta\log n$) and any promised norm bounds for $\|\mathbf{b}\|$ from the lattice analysis.
-21:if $\mathbf{A}\mathbf{b}\not\equiv\bm{0}\pmod{q}$ then
+21:if $\mathbf{U}^{\mathsf{T}}s+e\not\equiv\mathbf{t}\pmod{q}$ then
 22:  return fail
@@
 2:Output: $\mathbf{b}_{Q}=\mathbf{b}\bmod Q$, or fail.
-3:Let $U\in(\mathbb{Z}_{Q})^{N\times(n-1)}$ be columns $2..n$ of $\mathcal{S}$, and let $\bm{y}\in(\mathbb{Z}_{Q})^{N}$ be column 11 of $\mathcal{S}$.
-4:Solve $U\bm{z}\equiv\bm{y}\pmod{Q}$ and test uniqueness (e.g. via Smith normal form).
+3:Let $W\in(\mathbb{Z}_{Q})^{N\times(n-1)}$ be the submatrix of $\mathcal{S}$ consisting of columns $2..n$, and let $\bm{y}\in(\mathbb{Z}_{Q})^{N}$ be column 11 of $\mathcal{S}$.
+4:Solve $W\bm{z}\equiv\bm{y}\pmod{Q}$ and test uniqueness (e.g. via Smith normal form (kannan1979polynomial)).
 5:if no solution exists or the solution is not unique then
@@
 Remark 3.1 (Classical guessing loop for the norm).
-As in chen2024quantum, the small modulus is chosen as $Q=(c{+}1)\|\mathbf{b}\|^{2}$, where $\mathbf{b}$ is the (unknown) unique shortest vector of $L_{q}^{\perp}(\mathbf{A})$.
+As in chen2024quantum, the small modulus is chosen as $Q=(c{+}1)\|\mathbf{b}\|^{2}$, where $\mathbf{b}$ is the unknown unique shortest vector of $L_{q}^{\perp}(\mathbf{A})$.
 We run a classical outer loop over all polynomially many candidates $B_{\mathrm{guess}}$ for $\|\mathbf{b}\|^{2}$.
@@
 $\mathbf{b}=[-1,\,2p_{1}s^{\top},\,2p_{1}e^{\top}]^{\top}\in\mathbb{Z}^{n},$
-where $s$ is the LWE secret and $e$ is the error vector.
-Because $\|\mathbf{b}\|_{\infty}\leq 2p_{1}\beta\log n<Q/2$ in the parameter regime of interest, the reduction map $\mathbf{b}\mapsto\mathbf{b}\bmod Q$ is injective on the centred cube that contains the true shortest vector.
-Therefore, once Algorithm 1 recovers $\mathbf{b}\bmod Q$, the canonical lift to $(-Q/2,Q/2]^{n}$ recovers $\mathbf{b}$ as an integer vector.
-Dividing the last n1n-1 coordinates by $2p_{1}$ (after checking divisibility) yields $s$ and $e$, and the check $\mathbf{A}\mathbf{b}\equiv 0\pmod{q}$ certifies correctness.
+[…]
 Fourier sampling.
@@
 Proof of Lemma 3.3.
-Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals
+Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state Equation (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the unnormalized amplitude equals
 $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}\cdot\sum_{j\in\mathbb{Z}}\sum_{\bm{k}\in\mathcal{K}}\alpha_{E}(j,\bm{k})\,i^{\|\bm{k}\|^{2}}\,\omega_{M}^{\,j\langle\bm{u},\Delta\rangle+M_{2}\langle\bm{u},\bm{k}\rangle},$
-so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
+so the outcome probabilities are proportional to the squared magnitude of the inner sum.
 By AC5, conditioned on $E$ the coefficients factor as $\alpha_{E}(j,\bm{k})=\alpha_{E}(j)\beta_{E}(\bm{k})$.
@@
 where $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and we identify $t(\bm{u})$ with its representative in $\{0,\dots,Q-1\}$ when used inside the exponential.
+Equivalently, if we write $\bm{u}=\bm{u}_{Q}+Q\bm{t}$ with $\bm{u}_{Q}=\bm{u}\bmod Q$, then $\langle\bm{u},\Delta\rangle-\langle\bm{u}_{Q},\Delta\rangle=\langle Q\bm{t},2D^{2}\mathbf{b}\rangle\in M\mathbb{Z}$, so $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=\omega_{M}^{j\langle\bm{u}_{Q},\Delta\rangle}$ for all integers $j$.
 Therefore $S_{E}(\bm{u})$ factorizes as
-$S_{E}(\bm{u})=\Bigl(\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\,t(\bm{u})/Q}\Bigr)\cdot\Bigl(\sum_{\bm{k}\in\mathcal{K}}\beta_{E}(\bm{k})\,i^{\|\bm{k}\|^{2}}\,(-1)^{\langle\bm{u},\bm{k}\rangle}\Bigr)=\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)\cdot G_{E}(\bm{u}),$
-where $G_{E}(\bm{u})$ depends only on the parity pattern of $\bm{u}_{[2..n]}$. Thus
-$\Pr[\bm{u}]\ \propto\ |S_{E}(\bm{u})|^{2}=|\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)|^{2}\cdot|G_{E}(\bm{u})|^{2}.$
+$S_{E}(\bm{u})=\Bigl(\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\,t(\bm{u})/Q}\Bigr)\cdot\Bigl(\sum_{\bm{k}\in\mathcal{K}}\beta_{E}(\bm{k})\,i^{\|\bm{k}\|^{2}}\,(-1)^{\langle\bm{u},\bm{k}\rangle}\Bigr)=\widehat{\alpha}_{E}(t(\bm{u})/Q)\cdot G_{E}(\bm{u}),$
+where $G_{E}(\bm{u})$ depends only on the parity pattern of $\bm{u}_{[2..n]}$.
+For notational clarity, write $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and view $t(\bm{u})$ in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$.
+Then the envelope factor above is $\widehat{\alpha}_{E}(t(\bm{u})/Q)$.
+Thus
+$\Pr[\bm{u}]\ \propto\ |S_{E}(\bm{u})|^{2}=|\widehat{\alpha}_{E}(t(\bm{u})/Q)|^{2}\cdot|G_{E}(\bm{u})|^{2}.$
 For each residue $t\in\mathbb{Z}_{Q}$, let
@@
 Since $G_{E}(\bm{u})$ does not depend on $u_{1}$, the total grid mass $\sum_{\bm{u}\in T_{t}}|G_{E}(\bm{u})|^{2}$ is the same for every $t$.
-Hence the total probability mass of $T_{t}$ is proportional to $|\widehat{\alpha}_{E}(t/Q)|^{2}$, up to negligible truncation error.
-By the leakage bound (2.1) (good transcript) and this proportionality, the total off-resonant probability mass satisfies
+Hence, the total probability mass of $T_{t}$ is proportional to $|\widehat{\alpha}_{E}(t/Q)|^{2}$, up to negligible truncation error.
+By the leakage bound Equation (2.1) (good transcript) and this proportionality, the total off-resonant probability mass satisfies
 $\sum_{t\neq 0}\Pr[T_{t}]=\frac{\sum_{t\neq 0}|\widehat{\alpha}_{E}(t/Q)|^{2}}{\sum_{t\in\mathbb{Z}_{Q}}|\widehat{\alpha}_{E}(t/Q)|^{2}}\leq\operatorname{negl}(n),$
@@
 Assume AC1-AC5 and run Algorithm 2 in the correct iteration of the outer norm-guessing loop.
-Each sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ appended to $\mathcal{S}$ by Algorithm 1 is non-zero and satisfies
+Each accepted sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ appended to $\mathcal{S}$ by Algorithm 2 is non-zero by construction and satisfies
 $\langle\mathbf{b},\bm{u}_{Q}\rangle\equiv 0\pmod{Q}$
 except with probability $\operatorname{negl}(n)$ over the choice of a good transcript and the internal randomness of the measurements.
-Solving the resulting linear system modulo $Q$ (using the known normalization $b_{1}=-1$) recovers $\mathbf{b}\bmod Q$ and hence $\mathbf{b}$ as an integer vector.
+Solving the resulting linear system modulo $Q$ recovers $\mathbf{b}\bmod Q$ and hence $\mathbf{b}$ as an integer vector.
 The LWE secret and error are then extracted and verified in polynomial time with overwhelming probability.
@@
 Algorithm 2 makes only $\operatorname{poly}(n)$ oracle calls, so by a union bound we may condition on the event that all invocations are good, losing at most a negligible term in the overall success probability.
-In each oracle call within Algorithm 1, the sampling primitive Step9DaggerSample (Algorithm 2) measures $\tau=X_{1}\bmod D^{2}$ and applies the diagonal gate $U_{\mathrm{corr}}^{(\tau)}$.
-Lemma 2.4 (which uses the canonical-gauge requirement in AC4) shows that this removes the quadratic chirp and produces the chirp-free state $|\varphi_{7}\rangle$ assumed in AC5.
+In each oracle call within Algorithm 2, the sampling primitive Step9DaggerSample (Algorithm 1) measures $\tau=X_{1}\bmod D^{2}$, obtains the run-local residue $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from run-local side information (AC4), and applies the diagonal gate $U_{\mathrm{corr}}^{(v_{1,Q})}$.
+Lemma 2.4 shows that this removes the quadratic chirp and produces the chirp-free state $|\varphi_{7}\rangle$ assumed in AC5.
 Applying $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ and measuring yields $\bm{u}\in\mathbb{Z}_{M}^{n}$.
@@
 Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, the projection $H\to\mathbb{Z}_{Q}^{n-1}$ given by $\bm{u}\mapsto\bm{u}_{[2..n]}$ is a bijection.
-Hence the accepted vectors $\bm{u}_{Q,[2..n]}$ are i.i.d.
+Moreover, on $H$ the only vector with $\bm{u}_{[2..n]}=\bm{0}$ is $\bm{u}=\bm{0}$ (because $u_{1}\equiv 0$ then forces $u_{1}\equiv 0$), so rejecting $\bm{u}_{Q}=\bm{0}$ is equivalent to rejecting $\bm{u}_{Q,[2..n]}=\bm{0}$.
+Hence, the accepted vectors $\bm{u}_{Q,[2..n]}$ are i.i.d.
 (up to a negligible statistical error from the rare off-resonant outputs) uniform over $\mathbb{Z}_{Q}^{n-1}\setminus\{\bm{0}\}$.
-Let $U\in\mathbb{Z}_{Q}^{N\times(n-1)}$ be the matrix with these rows and let $\bm{y}$ be the vector of first coordinates.
-Then the true unknown $\mathbf{b}_{[2..n]}\bmod Q$ satisfies $U\mathbf{b}_{[2..n]}\equiv\bm{y}\pmod{Q}$.
-With overwhelming probability $U$ is injective as a map $(\mathbb{Z}_{Q})^{n-1}\to(\mathbb{Z}_{Q})^{N}$.
-Indeed, for any prime $p\mid Q$, the reduction $U\bmod p$ is (up to a negligible error from excluding the all-zero row) an i.i.d.
-uniform random matrix over the field $\mathbb{Z}_{p}$ with $N\geq C_{\mathrm{rank}}(n-1)$ rows, and standard rank bounds give $\Pr[\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)<n-1]\leq p^{-\Omega(n)}$.
-If $\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)=n-1$ and $U\bm{z}\equiv\bm{0}\pmod{p^{e}}$, then reducing modulo $p$ gives $\bm{z}\equiv\bm{0}\pmod{p}$; writing $\bm{z}=p\bm{z}_{1}$ and iterating shows $\bm{z}\equiv\bm{0}\pmod{p^{e}}$, so $\ker(U\bmod p^{e})=\{\bm{0}\}$.
-By the Chinese remainder theorem, $\ker(U\bmod Q)=\{\bm{0}\}$.
-Therefore the system has a unique solution $\bm{z}\in\mathbb{Z}_{Q}^{n-1}$, which Algorithm 4 computes (e.g.
-via Smith normal form) and uses to set $\mathbf{b}_{Q}=[-1,\bm{z}^{\top}]^{\top}=\mathbf{b}\bmod Q$.
+Let $W\in\mathbb{Z}_{Q}^{N\times(n-1)}$ be the matrix with these rows and let $\bm{y}$ be the vector of first coordinates.
+Then the true unknown $\mathbf{b}_{[2..n]}\bmod Q$ satisfies $W\mathbf{b}_{[2..n]}\equiv\bm{y}\pmod{Q}$.
+With overwhelming probability $W$ is injective as a map $(\mathbb{Z}_{Q})^{n-1}\to(\mathbb{Z}_{Q})^{N}$.
+Because $Q=\operatorname{poly}(n)$, the number of distinct prime divisors of $Q$ is at most $\log_{2}Q=O(\log n)$, so a union bound over $p\mid Q$ preserves a negligible overall failure probability.
+Indeed, for any prime $p\mid Q$, the reduction $W\bmod p$ is (up to a negligible error from excluding the all-zero row) an i.i.d.
+uniform random matrix over the field $\mathbb{Z}_{p}$ with $N\geq C_{\mathrm{rank}}(n-1)$ rows, and standard rank bounds give $\Pr[\operatorname{rank}_{\mathbb{Z}_{p}}(W\bmod p)<n-1]\leq p^{-\Omega(n)}$.
+If $\operatorname{rank}_{\mathbb{Z}_{p}}(W\bmod p)=n-1$ and $W\bm{z}\equiv\bm{0}\pmod{p^{e}}$, then reducing modulo $p$ gives $\bm{z}\equiv\bm{0}\pmod{p}$; writing $\bm{z}=p\bm{z}_{1}$ and iterating shows $\bm{z}\equiv\bm{0}\pmod{p^{e}}$, so $\ker(W\bmod p^{e})=\{\bm{0}\}$.
+By the Chinese remainder theorem, $\ker(W\bmod Q)=\{\bm{0}\}$.
+Therefore the system has a unique solution $\bm{z}\in\mathbb{Z}_{Q}^{n-1}$, which Algorithm 4 computes and uses to set $\mathbf{b}_{Q}=[-1,\bm{z}^{\top}]^{\top}=\mathbf{b}\bmod Q$.
 Finally, the bound $\|\mathbf{b}\|_{\infty}<Q/2$ from (R1) makes the centered lift from $\mathbb{Z}_{Q}^{n}$ to $(-Q/2,Q/2]^{n}$ injective on the relevant region, so lifting $\mathbf{b}_{Q}$ recovers $\mathbf{b}$ as an integer vector.
-The subsequent divisibility, planted-prefix, shortness, and lattice-membership checks (Algorithm 3) succeed for the correct guess and reject incorrect guesses, as in chen2024quantum.
+The subsequent divisibility, planted-prefix, shortness, and LWE-consistency checks (Algorithm 3) succeed for the correct guess and reject incorrect guesses, as in chen2024quantum.
 ∎
@@
 Via the classical reductions of chen2024quantum, this yields a quantum algorithm for standard LWE in the same parameter regime, conditional on AC1-AC5.
+4 Related Work
+Lattice-based cryptography is rooted in worst-case hardness frameworks such as Ajtai’s construction of hard lattice instances (ajtai1996generating) and subsequent worst-case-to-average-case reductions based on Gaussian measure (micciancio2007worst).
+Regev introduced LWE and gave quantum worst-case reductions from lattice problems such as GapSVP and SIVP to LWE (regev2009lattices).
+Subsequent work clarified parameter tradeoffs and established classical hardness in a range of regimes (peikert2009public; brakerski2013classical).
+These reductions underpin a broad ecosystem of cryptographic constructions, including trapdoors for hard lattices (gentry2008trapdoors) and fully homomorphic encryption (gentry2009fully).
+For context, the fastest known worst-case SVP algorithms are classical and typically rely on sieving and/or discrete Gaussians, starting from the sieve of Ajtai-Kumar-Sivakumar (ajtai2001sieve) and later practical refinements (nguyen2008sieve), and including the $2^{n+o(n)}$-time discrete-Gaussian algorithm of aggarwal2015solving.
+Discrete Gaussian sampling is also a central technical tool in lattice algorithms more broadly; efficient lattice Gaussian samplers were developed, for example, by Peikert (peikert2010efficient).
+On the quantum side, lattice algorithms are often organized around preparing structured coset- or line-supported superpositions and extracting information via Fourier sampling.
+A canonical early connection is Regev’s reduction from unique-SVP to the dihedral hidden subgroup problem (DHSP) via coset sampling (regev2004quantum), together with subexponential-time algorithms for DHSP such as Kuperberg’s sieve (kuperberg2005subexponential).
+Related dihedral-coset models have since appeared in lattice settings, e.g., the extrapolated dihedral coset problem (EDCP), which is closely connected to LWE (brakerski2018learning).
+On the average-case state side, Chen, Liu, and Zhandry introduced a filtering framework for decoding hidden linear structure from LWE-like quantum states and used it to obtain polynomial-time quantum algorithms for several average-case variants, including EDCP in certain regimes (chen2022quantum).
+Subsequent work has studied both algorithms and hardness barriers for quantum LWE-state families with structured (e.g., Gaussian) amplitudes and linear/quadratic phase terms (chen2025lwe).
+Chen’s Karst-wave construction (chen2024quantum) is an instance of this Fourier-sampling template with amplitude engineering via complex-Gaussian windowing, yielding a post-processing state supported on a shifted grid line whose amplitudes include a quadratic chirp in the loop index.
+Chen’s published post-processing (Steps 8-9) exploits additional arithmetic structure to handle the unknown shift and chirp.
+The present work isolates an explicit access-model promise that suffices for exact post-processing in the LWE regime: run-local access to $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (AC4), enabling single-coordinate chirp cancellation followed by a direct $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ without learning the full offset $\bm{v}^{\ast}(E)$.
 5 Conclusion
@@
 Under these assumptions, we obtain a quantum algorithm that solves the chosen-secret LWE problem of chen2024quantum in almost the same parameter regime as in that work.
-Acknowledgment
+Acknowledgments
 We are grateful to all who provided constructive discussions and helpful feedback.
+The author used AI-enabled tools solely for English grammar and clarity suggestions in non-technical prose.
+All technical ideas, proofs, and results are the author’s own work.